### v1

# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

Abstract
When estimating the direct and indirect causal effects using the influence function-based estimator of the mediation functional, it is crucial to understand what aspects of the treatment, the mediator, and the outcome mean mechanisms should be focused on. Specifically, considering them as nuisance functions and attempting to fit these nuisance functions as accurate as possible is not necessarily the best approach to take. In this work, we propose a two-stage estimation strategy for the nuisance functions that estimates the nuisance functions based on the role they play in the structure of the bias of the influence function-based estimator of the mediation functional. We provide robustness analysis of the proposed method, as well as sufficient conditions for consistency and asymptotic normality of the estimator of the parameter of interest.
Keywords: Causal Mediation Analysis; Influence Function-Based Estimation; Multiple Robustness; Balancing Estimators.
1 Introduction
In contrast to simply assessing the overall treatment effect of a treatment variable on an outcome variable of interest, causal mediation analysis delves deeper by posing a “how” inquiry, investigating the underlying mechanisms through which the treatment influences the outcome. Due to this conceptually invaluable focus, it has gained significant popularity in many fields of science including social science, economics, and epidemiology (VanderWeele,, 2015; MacKinnon,, 2012; Hayes,, 2017; Richiardi et al.,, 2013; Ten Have and Joffe,, 2012).
The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model for the data generating process, the direct and indirect effects can be identified by a functional known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those assumptions hold in the system under study and the parameter is identified, the next question is how to estimate the mediation functional.
There has been several estimation strategies proposed in the literature based on regression and weighting techniques. In their seminal paper, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric effeicint and has the desired property that under certain regularity conditions, it remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism, one describing the mediator mechanism, and one describing the outcome mean mechanism, and the model for two out of three should be correctly specified. Among the aforementioned nuisance functions, the mediator mechanism, that is, the conditional distribution of the mediator variable given covariates and the treatment, is usually a difficult object to estimate. Trying to naively fit this conditional distribution as good as possible will most likely lead to a final estimator of the mediation functional that is sensitive to even small errors in the estimator of this nuisance function. More importantly, in the process of estimation, it is not clear what aspects of this object should be focused on to obtain a better estimate of the mediation functional.
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis, which avoids direct estimation of the mediator mechanism. In the first stage, we estimate two basic nuisance functions, namely the treatment mechanism and the outcome mean mechanism. In the second stage, using the basic nuisance functions, we estimate two additional nuisance functions that contain the needed information regarding the mediator mechanism. We use the weighted balancing approach of Imai and Ratkovic, (2014) to design the estimator of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide parametric and nonparametric versions of the balancing estimators. The other two nuisance functions are obtained using standard parametric or nonparametric regressions. After obtaining all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We provide a robustness analysis of the resulting estimator and demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain a consistent estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest.
The rest of the paper is organized as follows. In Section 2, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. In Section 3, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. We propose our two-stage nuisance function estimation approach in Section 4, and we provide the robustness and asymptotic analyses of the proposed estimator in Section 5. We conclude in Section 6.
2 Model Description
$A$$X$$M$$Y$
Figure 1: A graphical causal model for the mediation model.
We consider a model with binary treatment variable $A\in\{0,1\}$, observed pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$, and $Y^{(m)}=Y^{(A,m)}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $10<p(m\mid a,x)<$ 1, and $10<p(a\mid x)<$ 1.
We also require that the model does not contain any unobserved confounders, which is a common and standard assumption in the causal mediation analysis literature. This can be formalized using the sequential exchangeability assumption (Imai et al.,, 2010).
Assumption 1 (Sequential exchangeability).
For any $a,a^{\prime}\in\{0,1\}$, and any value of the mediator $m$, we have
  1. (i)
    $Y^{(a,m)}\perp\mkern-9.5mu\perp A\mid X$,
  2. (ii)
    $M^{(a)}\perp\mkern-9.5mu\perp A\mid X$,
  3. (iii)
    $Y^{(a,m)}\perp\mkern-9.5mu\perp M^{(a^{\prime})}\mid X$.
Intuitively, Part $(i)$ of the assumption requires that the treatment-outcome and mediator-outcome relations are not confounded by unobserved variables, Part $(ii)$ of the assumption requires that the treatment-mediator relation is not confounded by unobserved variables (see Figure 1), and Part $(iii)$ of the assumption, which is a so-called cross-world independence assumption, requires that the potential outcomes of $Y$ and $M$ even for different values of the treatment are conditionally independent. All three parts are untestable and should be decided based on expert’s knowledge and the researcher’s belief about how the variables in the system are generated.
Consider the parameter average treatment effect (ATE) of the treatment $A$ on the outcome $Y$, represented by $\theta_{ATE}=\mathbb{E}[Y^{(1)}-Y^{(0)}]$. The main goal in causal mediation analysis is to quantify the part of the ATE that is mediated by the variable $M$, and the part that is directly reaching the outcome variable. To this end, the ATE can be partitioned as follows (Robins and Greenland,, 1992; Pearl,, 2001).
$\theta_{ATE}$ $=\mathbb{E}[Y^{(1)}-Y^{(0)}]$
$=\mathbb{E}[Y^{(1,M^{(1)})}-Y^{(0,M^{(0)})}]$
$=\underbrace{\mathbb{E}[Y^{(1,M^{(1)})}-Y^{(1,M^{(0)})}]}_{\text{NIE}}+\underbrace{\mathbb{E}[Y^{(1,M^{(0)})}-Y^{(0,M^{(0)})}]}_{\text{NDE}}.$
The first and the second terms in the last expression are called the natural indirect effect (NIE) and the natural direct effect (NDE) of the treatment on the outcome, respectively (Pearl,, 2001). [^1] These two terms are called the total indirect effect and the pure direct effect, respectively by Robins and Greenland, (1992). NIE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is fixed at $a=1$, while the mediator behaves as if the treatment had been changed from value $a=0$ to $a=1$. NDE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is changed from value $a=0$ to $a=1$, while the mediator behaves as if the treatment is fixed at $a=0$.
We notice that by consistency, positivity, and sequential exchangeability, we have
$\mathbb{E}[Y^{(1,M^{(1)})}]=\mathbb{E}[Y^{(1)}]=\mathbb{E}[Y^{(1)}\mid A=1]=\mathbb{E}[Y\mid A=1].$
That is $\mathbb{E}[Y^{(1,M^{(1)})}]$ is identified. Similar argument holds for $\mathbb{E}[Y^{(0,M^{(0)})}]$. The parameter $\theta_{0}=\mathbb{E}[Y^{(1,M^{(0)})}]$ requires more caution. It is known that under the assumptions of our model, this parameter can be identified as follows (Pearl,, 2001; Imai et al.,, 2010).
$\theta_{0}$ $=\mathbb{E}[Y^{(1,M^{(0)})}]$
[…]
$=:\psi_{0}.$
The last expression is called the mediation functional (also known as mediation formula). Note that this expression is purely a functional of the observational data. We denote the mediation functional by $\psi_{0}$. Note that under the identification assumptions in our model, $\psi_{0}=\theta_{0}$, but in general, these two parameters are not necessarily equal. In the sequel, we refer to $\psi_{0}$ as the parameter of interest.
3 Influence Function-Based Estimator
Tchetgen Tchetgen and Shpitser, (2012) showed that the efficienet influence function of the mediation functional is given by
$IF_{\psi_{0}}=$ $\frac{I(A=1)p(M\mid A=0,X)}{p(A=1\mid X)p(M\mid A=1,X)}\{Y-\mathbb{E}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{p(A=0\mid X)}\{\mathbb{E}[Y\mid A=1,M,X]-\eta(1,0,X)\}+\eta(1,0,X)-\psi_{0},$
where
[…]
Based on $IF_{\psi_{0}}$, they proposed to first estimate the following nuisance functions.
  • $p(A=1\mid X)$, $\forall X\in\mathcal{X}$,
  • $\mathbb{E}[Y\mid A=1,M,X]$, $\forall M\in\mathcal{M},X\in\mathcal{X}$,
  • $p(M\mid A=a,X)$, $\forall M\in\mathcal{M},X\in\mathcal{X}$.
Then, estimate the mediation functional by
$\hat{\psi}^{TTS}=$ $\mathbb{E}_{n}\bigg{[}\frac{I(A=1)\hat{p}(M\mid A=0,X)}{\hat{p}(A=1\mid X)\hat{p}(M\mid A=1,X)}\{Y-\hat{\mathbb{E}}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{1-\hat{p}(A=1\mid X)}\{\hat{\mathbb{E}}[Y\mid A=1,M,X]-\hat{\eta}(1,0,X)\}+\hat{\eta}(1,0,X)\bigg{]},$
where $\mathbb{E}_{n}$ is the empirical expectation operator (treating nuisance functions as non-random), and
[…]
The authors showed that this estimator is multiply robust in the following sense.
Proposition 1.
$\hat{\psi}^{TTS}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  • $\{\hat{p}(A=1\mid X),\hat{\mathbb{E}}[Y\mid A=1,M,X]\}$,
  • $\{\hat{p}(A=1\mid X),\hat{p}(M\mid A=a,X)\}$,
  • $\{\hat{\mathbb{E}}[Y\mid A=1,M,X],\hat{p}(M\mid A=a,X)\}$.
4 Proposed Estimation Approach
We define the following nuisance functions.
$\pi_{1}(X)=p(A=0\mid X),$ $\forall X\in\mathcal{X},$
$\pi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M\mid A=1,X)},$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
$\mu_{1}(M,X)=\mathbb{E}[Y\mid A=1,M,X],$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
$\mu_{2}(X)=\mathbb{E}[\mu_{1}(M,X)\mid A=0,X],$ $\forall X\in\mathcal{X}.$
Note that the parameter of interest can be written as
$\psi_{0}=\mathbb{E}[\mu_{2}(X)].$
We start by presenting a high-level view of our proposed approach.
Estimating the nuisance functions: In order for estimating the nuisance functions, we proceed in the following two stages.
  • Stage 1:
    • We estimate $\pi_{1}$ using the following balancing estimator:
      We choose $\hat{\pi}_{1}$ such that it satisfies
      $\mathbb{E}\left[\left\{\frac{1-A}{\hat{\pi}_{1}(X)}-1\right\}h_{1}(X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1},$ (1)
      where $\mathcal{H}_{1}\subset L^{2}(X)$ is a function space chosen by the researcher.
    • We estimate $\mu_{1}$ by regressing $Y$ on $\{M,X,A=1\}$.
  • Stage 2:
    • We estimate $\pi_{2}$ using the following balancing estimator:
      Given $\hat{\pi}_{1}$ from Stage 1, we choose $\hat{\pi}_{2}$ such that it satisfies
      $\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{1-A}{\hat{\pi}_{1}(X)}\right\}h_{2}(M,X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2},$ (2)
      where $\mathcal{H}_{2}\subset L^{2}(M,X)$ is a function space chosen by the researcher.
    • Given $\hat{\mu}_{1}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1}$ on $\{X,A=0\}$.
Estimating the parameter of interest: Our final estimator of the parameter of interest will be the following.
$\hat{\psi}^{mr}=\mathbb{E}_{n}\left[\hat{\mu}_{2}(X)+\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}\{Y-\hat{\mu}_{1}(M,X)\}\right].$
In Stage 1, we can choose to use a parametric model $\{\pi_{1}(X;\alpha):\alpha\in\Xi_{\alpha}\}$, parametrized by $\alpha$ for $\pi_{1}$. Suppose $\Xi_{\alpha}=\mathbb{R}^{d}$. Then one can choose $d$ score functions $\{h_{1,1},...,h_{1,d}\}$ from $\mathcal{H}_{1}$, and obtain $\hat{\alpha}$ using generalized method of moments (GMM) by solving the following system of $d$ equations.
$\begin{cases}&\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X;\hat{\alpha})}-1\right\}h_{1,1}(X)\right]=0\\ &\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X;\hat{\alpha})}-1\right\}h_{1,2}(X)\right]=0\\ &\vdots\\ &\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X;\hat{\alpha})}-1\right\}h_{1,d}(X)\right]=0\end{cases}$
Alternatively, one can choose to proceed with a nonparametric estimator. A naive approach would be by solving the following optimization problem.
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)\right],$
where $\Pi_{1}\subset L^{2}(X)$ is a function space chosen by the researcher. Unfortunately, this estimator will be unstable and highly sensitive to noise. The following result suggests an alternative.
Proposition 2.
$\hat{\pi}_{1}$ obtained from the following minimax optimization satisfies equation (1) with equality.
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(X)}\max_{h_{1}\in L^{2}(X)}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
Inspired by Proposition 2, we estimate $\pi_{1}$ as
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$ (3)
Note that our proposed nonparametric estimator $\hat{\pi}_{1}$ in equation (3) can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect.
Similarly, in Stage 2, to estimate $\pi_{2}$, we can use a GMM estimator, or use the following counterpart of estimator (3).
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{1-A}{\hat{\pi}_{1}(X)}\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right],$ (4)
where $\Pi_{2}\subset L^{2}(M,X)$ is a function space chosen by the researcher.
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, any standard parametric or nonparametric regression method can be used.
5 Analyses
We start by making the observation that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions in (1) and (2) with equality.
Proposition 3.
  1. (a)
    If $\hat{\pi}_{1}=\pi_{1}$, then
    $\mathbb{E}\left[\left\{\frac{1-A}{\hat{\pi}_{1}(X)}-1\right\}h_{1}(X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1}.$
  2. (b)
    If $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$, then
    $\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{1-A}{\hat{\pi}_{1}(X)}\right\}h_{2}(M,X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2}.$
Suppose conditions in (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(X)$ and $\mathcal{H}_{2}=L^{2}(M,X)$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are not the corresponding $L^{2}$ spaces, we can have conditions in (1) and (2) hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions in (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{mr}$.
Theorem 1.
The bias of the estimator $\hat{\psi}^{mr}$ satisfies the following equality.
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$ $=\mathbb{E}\left[\left\{\frac{(1-A)}{\hat{\pi}_{1}(X)}-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
From the structure of the bias of the estimator $\hat{\psi}^{mr}$, it is evident that if $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$ can be well-approximated by the functions in $\mathcal{H}_{1}$, and $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$ can be well-approximated by the functions in $\mathcal{H}_{2}$, then we can eliminate the bias of the estimator $\hat{\psi}^{mr}$ even if we have not had a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
Importantly, our approach retains the multiple-robustness property, as stated in the following result.
Proposition 4.
$\hat{\psi}^{mr}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  • $\{\hat{\pi}_{1},\hat{\pi}_{2}\}$
  • $\{\hat{\mu}_{1},\hat{\mu}_{2}\}$
  • $\{\hat{\pi}_{1},\hat{\mu}_{1}\}$
Proposition 4 immediately follows from Proposition 3 and Theorem 1. Figure 2 summarizes the result in Proposition 4.
$\hat{\pi}_{1}$$\hat{\pi}_{2}$$\hat{\mu}_{1}$$\hat{\mu}_{2}$tage 1:Stage 2:
Figure 2: Nuisance functions estimated in each stage of the approach. The arrow from $\hat{\pi}_{1}$ to $\hat{\pi}_{2}$ indicates that the estimator of $\pi_{1}$ is used in estimating $\pi_{2}$. The arrow from $\hat{\mu}_{1}$ to $\hat{\mu}_{2}$ is interpreted similarly. Correct specification of any of the pairs of functions in an ellipse suffices unbiasedness of the final estimator of the parameter of interest.
5.1 Asymptotic Normality
We next provide sufficient conditions for consistency and asymptotic normality of the estimator $\hat{\psi}^{mr}$.
Assumption 2.
  1. (a)
    $\big{\|}\frac{p(A=0\mid X)}{\hat{\pi}_{1}(X)}-1\big{\|}_{2}=o_{p}(1)$, $\big{\|}\frac{p(A=1\mid M,X)\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{p(A=0\mid M,X)}{\hat{\pi}_{1}(X)}\big{\|}_{2}=o_{p}(1)$,
    $\|\mu_{2}(X)-\hat{\mu}_{2}(X)\|_{2}=o_{p}(1)$, and $\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\|_{2}=o_{p}(1)$.
  2. (b)
    $\big{\|}\frac{p(A=0\mid X)}{\hat{\pi}_{1}(X)}-1\big{\|}_{2}\|\mu_{2}(X)-\hat{\mu}_{2}(X)\|_{2}=o_{p}(n^{-1/2})$, and
    $\big{\|}\frac{p(A=1\mid M,X)\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{p(A=0\mid M,X)}{\hat{\pi}_{1}(X)}\big{\|}_{2}\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\|_{2}=o_{p}(n^{-1/2})$.
Define $O:=(X,A,M,Y)$, and for any estimates of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
$m(O;\tilde{\pi}_{1}$ $,\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}):=\tilde{\mu}_{2}(X)+\frac{(1-A)}{\tilde{\pi}_{1}(X)}\{\tilde{\mu}_{1}(M,X)-\tilde{\mu}_{2}(X)\}+\frac{A\tilde{\pi}_{2}(M,X)}{1-\tilde{\pi}_{1}(X)}\{Y-\tilde{\mu}_{1}(M,X)\}.$
We have the following result regarding the asymptotic normality of the estimator $\hat{\psi}^{mr}$.
Theorem 2.
Suppose nuisance functions and the parameter of interest are estimated on disjoint samples with sample sizes of the same order as the total sample size. Under Assumption 2, the estimator $\hat{\psi}^{mr}$ is asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})=\sqrt{n}\mathbb{E}_{n}\left[m(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]+o_{p}(1).$
As a corollary of Theorem 2, one can use the influence function of the proposed estimator to obtain confidence intervals for the parameter of interest.
6 Conclusion
In estimating a functional of interest that involves nuisance functions, if one naively aims to perfectly fit the nuisance functions, the resulting estimator of the parameter of interest can be sensitive to even small errors in the nuisance function estimators. Therefore, in order to estimate a functional, it is important to investigate the role of nuisance functions involved in that functional in the structure of the bias. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional when the influence function-based estimator is used for the parameter of interest. Our proposed estimators were based on the approach of balancing that can target the bias of the parameter of interest more directly and also can lead to an estimator which is less sensitive to error. We demonstrated that this approach retains the multiple-robustness property. Moreover, we provided conditions on our nuisance function estimators sufficient for asymptotic normality of the influence function-based estimator of the mediation functional.
Appendix A Proofs
Proof of Proposition 2.
We note that
$\arg\max_{h_{1}}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left(\frac{1}{2}h_{1}(X)-\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}.$
Therefore,
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[2\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}-\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[\mathbb{E}\bigg{[}\frac{1-A}{\pi_{1}(X)}-1\bigg{|}X\bigg{]}^{2}\right],$
which since zero is achievable by $p(A=0\mid X)$, it implies that
$\mathbb{E}\bigg{[}\frac{1-A}{\hat{\pi}_{1}(X)}-1\bigg{|}X\bigg{]}=0.$
Therefore, for all $h_{1}\in\mathcal{H}_{1}$,
$\mathbb{E}\left[\left\{\frac{1-A}{\hat{\pi}_{1}(X)}-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\mathbb{E}\bigg{[}\frac{1-A}{\hat{\pi}_{1}(X)}-1\bigg{|}X\bigg{]}h_{1}(X)\right]=0.$
Proof of Proposition 3.
  1. (a)
    For all $h_{1}\in\mathcal{H}_{1}$, we have
    $\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\left\{\frac{\mathbb{E}[1-A\mid X]}{\pi_{1}(X)}-1\right\}h_{1}(X)\right]=0.$
  2. (b)
    For all $h_{2}\in\mathcal{H}_{2}$, we have
    $\mathbb{E}\left[\left\{\frac{A\pi_{2}(M,X)}{1-\pi_{1}(X)}-\frac{1-A}{\pi_{1}(X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(A=1\mid M,X)\frac{p(M\mid A=0,X)}{p(M\mid A=1,X)}}{p(A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(M\mid A=0,X)}{p(M\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}-\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=0$
Proof of Theorem 1.
$\mathbb{E}[\hat{\theta}^{mr}]-\theta_{0}$
$=\mathbb{E}\left[\hat{\mu}_{2}(X)+\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}\{Y-\hat{\mu}_{1}(M,X)\}\right]-\mathbb{E}[\mu_{2}(X)]$
$=\mathbb{E}\left[\hat{\mu}_{2}(X)-\mu_{2}(X)+\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}\{Y-\hat{\mu}_{1}(M,X)\}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}\{Y-\hat{\mu}_{1}(M,X)\}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\bigg{|}X\right]-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\mathbb{E}\left[\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}\{Y-\hat{\mu}_{1}(M,X)\}\bigg{|}M,X\right]-\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}[(1-A)\mu_{1}(M,X)\mid X]-\pi_{1}(X)\hat{\mu}_{2}(X)}{\hat{\pi}_{1}(X)}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\hat{\pi}_{2}(M,X)\frac{\mathbb{E}[AY\mid M,X]-\mathbb{E}[A\mid M,X]\hat{\mu}_{1}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}[(1-A)\mid X]\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}}{\hat{\pi}_{1}(X)}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\frac{\hat{\pi}_{2}(M,X)\mathbb{E}[A\mid M,X]\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}}{1-\hat{\pi}_{1}(X)}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{\frac{(1-A)}{\hat{\pi}_{1}(X)}-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{(1-A)}{\hat{\pi}_{1}(X)}\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$