### v2

# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

Abstract
Tchetgen Tchetgen and Shpitser (2012) introduced an efficient, debiased, and robust influence function-based estimator for the mediation functional, which is the key component in mediation analysis. This estimator relies on the treatment, mediator, and outcome mean mechanisms. However, treating these three mechanisms as nuisance functions and fitting them as accurately as possible may not be the most effective approach. Instead, it is essential to identify the specific functionals and aspects of these mechanisms that impact the estimation of the mediation functional. In this work, we propose a two-stage estimation strategy for certain nuisance functions in the influence function of the mediation functional that are based on these three mechanisms. This strategy is guided by the role those nuisance functions play in the bias structure of the influence function-based estimator for the mediation functional. In the first stage, we estimate two primary nuisance functions, namely the inverse treatment mechanism and the outcome mean mechanism. In the second stage, we leverage these primary functions to estimate two additional nuisance functions that encapsulate the needed information about the mediator mechanism. We propose a nonparametric weighted balancing estimation approach to design the estimator for one of the nuisance functions in Stage 1 and one of in Stage 2, where the weights are designed directly based on the bias of the final estimator for the mediation functional. The remaining two nuisance functions are estimated using standard parametric or nonparametric regression methods. Once all four nuisance functions are obtained, they are incorporated into the influence function-based estimator. We provide a robustness analysis of the proposed method and establish sufficient conditions for consistency and asymptotic normality of our estimator for the mediation functional.
Keywords: Causal Mediation Analysis; Influence Function-Based Estimation; Multiple Robustness; Balancing Estimators.
1 Introduction
In contrast to simply assessing the overall treatment effect of a treatment variable on an outcome variable of interest, causal mediation analysis delves deeper by posing a “how” inquiry, investigating the underlying mechanisms through which the treatment influences the outcome. Due to this conceptually invaluable focus, it has gained significant popularity in many fields of science including social science, economics, and epidemiology (VanderWeele,, 2015; MacKinnon,, 2012; Hayes,, 2017; Richiardi et al.,, 2013; Ten Have and Joffe,, 2012).
The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model for the data generating process, the direct and indirect effects can be identified by a functional known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those assumptions hold in the setting under study and the parameter is identified, the next question is how to estimate the mediation functional.
There has been several estimation strategies proposed in the literature based on regression and weighting techniques. In their seminal paper, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric efficient and has the desired property that under certain regularity conditions, it remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism (i.e., the conditional distribution of the treatment given the covariates), one describing the mediator mechanism (i.e., the conditional distribution of the mediator variable given covariates and the treatment), and one describing the outcome mean mechanism (i.e., the conditional expectation of the outcome variable given covariates, the treatment, and the mediator), and the model for two out of three should be correctly specified. There are two challenges with their proposed estimator: (1) among the aforementioned nuisance functions, the mediator mechanism is usually a difficult object to estimate, as it may be the density of a multivariate continuous variable conditioned on another multivariate continuous variable; (2) in their estimator, the estimator of the treatment and mediator mechanisms appear in the denominator, and hence can significantly inflate the variance of the estimator, especially when these quantities are small. Therefore, trying to naively fit these conditional distributions as good as possible will not necessarily lead to an estimator with reasonable mean squared error. Instead, it is important to understand what aspects of these objects should be focused on to obtain a better estimate of the mediation functional.
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis, which avoids direct estimation of the treatment and mediator mechanisms. In the first stage, we estimate two basic nuisance functions, namely the inverse of the treatment mechanism and the outcome mean mechanism. In the second stage, using the basic nuisance functions, we estimate two additional nuisance functions that contain the needed information regarding the mediator mechanism. We propose a non-parametric version of the weighted balancing approach of Imai and Ratkovic, (2014) to design the estimator of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide an implementation of our non-parametric estimators with RKHSes as the function classes, and provide closed-form solutions for the nuisance parameters. The other two nuisance functions are obtained using standard parametric or nonparametric regressions. After obtaining all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We provide a robustness analysis of the resulting estimator and demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain a asymptotically unbiased estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest.
The rest of the paper is organized as follows. In Section 2, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. In Section 3, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. We propose our two-stage nuisance function estimation approach as well as its implementation using RKHSes in Section 4. We provide the robustness and asymptotic analyses of the proposed estimator in Section 5. We conclude in Section 6.
2 Model Description
$A$$X$$M$$Y$
Figure 1: A graphical causal model for the mediation model.
We consider a model with binary treatment variable $A\in\{0,1\}$, pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. Let $P_{0}$ represent the true underlying distribution from which the independent and identically distributed observations $\{O_{i}:=(X,A,M,Y)_{i}\}_{i=1}^{n}$ are collected. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$, and $Y^{(m)}=Y^{(A,m)}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $p(m\mid a,x)>0$, and $p(a\mid x)>0$.
We also require that the model does not contain any unobserved confounders, which is a common and standard assumption in the causal mediation analysis literature. This can be formalized using the sequential exchangeability assumption (Imai et al.,, 2010).
Assumption 1 (Sequential exchangeability).
For any $a,a^{\prime}\in\{0,1\}$, and any value of the mediator $m$, we have
  1. (i)
    $Y^{(a,m)}\perp\mkern-9.5mu\perp\{A,M\}\mid X$,
  2. (ii)
    $M^{(a)}\perp\mkern-9.5mu\perp A\mid X$,
  3. (iii)
    $Y^{(a,m)}\perp\mkern-9.5mu\perp M^{(a^{\prime})}\mid X$.
Intuitively, Part $(i)$ of the assumption requires that the treatment-outcome and mediator-outcome relations are not confounded by unobserved variables, Part $(ii)$ of the assumption requires that the treatment-mediator relation is not confounded by unobserved variables (see Figure 1), and Part $(iii)$ of the assumption, which is a so-called cross-world independence assumption, requires that the potential outcomes of $Y$ and $M$ even for different values of the treatment are conditionally independent. All three parts are untestable and should be decided based on expert’s knowledge and the researcher’s belief about how the variables in the system are generated.
Consider the parameter average treatment effect (ATE) of the treatment $A$ on the outcome $Y$, represented by $\theta_{ATE}=\mathbb{E}[Y^{(1)}-Y^{(0)}]$. The main goal in causal mediation analysis is to quantify the part of the ATE that is mediated by the variable $M$, and the part that is directly reaching the outcome variable. To this end, the ATE can be partitioned as follows (Robins and Greenland,, 1992; Pearl,, 2001).
$\theta_{ATE}$ $=\mathbb{E}[Y^{(1)}-Y^{(0)}]$
$=\mathbb{E}[Y^{(1,M^{(1)})}-Y^{(0,M^{(0)})}]$
$=\underbrace{\mathbb{E}[Y^{(1,M^{(1)})}-Y^{(1,M^{(0)})}]}_{\text{NIE}}+\underbrace{\mathbb{E}[Y^{(1,M^{(0)})}-Y^{(0,M^{(0)})}]}_{\text{NDE}}.$
The first and the second terms in the last expression are called the natural indirect effect (NIE) and the natural direct effect (NDE) of the treatment on the outcome, respectively (Pearl,, 2001). [^1] These two terms are called the total indirect effect and the pure direct effect, respectively by Robins and Greenland, (1992). NIE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is fixed at $a=1$, while the mediator behaves as if the treatment had been changed from value $a=0$ to $a=1$. NDE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is changed from value $a=0$ to $a=1$, while the mediator behaves as if the treatment is fixed at $a=0$.
We notice that by consistency, positivity, and sequential exchangeability, we have
$\mathbb{E}[Y^{(1,M^{(1)})}]=\mathbb{E}[Y^{(1)}]=\mathbb{E}[\mathbb{E}[Y^{(1)}\mid X]]=\mathbb{E}[\mathbb{E}[Y^{(1)}\mid A=1,X]]=\mathbb{E}[\mathbb{E}[Y\mid A=1,X]].$
That is $\mathbb{E}[Y^{(1,M^{(1)})}]$ is identified. Similar argument holds for $\mathbb{E}[Y^{(0,M^{(0)})}]$. The parameter $\theta_{0}=\mathbb{E}[Y^{(1,M^{(0)})}]$ requires more caution. It is known that under the assumptions of our model, this parameter can be identified as follows (Pearl,, 2001; Imai et al.,, 2010).
$\theta_{0}$ $=\mathbb{E}[Y^{(1,M^{(0)})}]$
[…]
$=:\psi_{0}.$
The last expression is called the mediation functional (also known as mediation formula). Note that this expression is purely a functional of the observational data. We denote the mediation functional by $\psi_{0}$. Note that under the identification assumptions in our model, $\psi_{0}=\theta_{0}$, but in general, these two parameters are not necessarily equal. In the sequel, we refer to $\psi_{0}$ as the parameter of interest.
3 Influence Function-Based Estimator
Tchetgen Tchetgen and Shpitser, (2012) showed that in a nonparametric model, the efficient influence function of the mediation functional is given by
$IF_{\psi_{0}}=$ $\frac{I(A=1)p(M\mid A=0,X)}{p(A=1\mid X)p(M\mid A=1,X)}\{Y-\mathbb{E}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{p(A=0\mid X)}\{\mathbb{E}[Y\mid A=1,M,X]-\eta(1,0,X)\}+\eta(1,0,X)-\psi_{0},$
where
[…]
Based on $IF_{\psi_{0}}$, they proposed to first estimate the following nuisance functions.
  • $p(A=1\mid X)$, $\forall X\in\mathcal{X}$,
  • $\mathbb{E}[Y\mid A=1,M,X]$, $\forall M\in\mathcal{M},X\in\mathcal{X}$,
  • $p(M\mid A=a,X)$, $\forall M\in\mathcal{M},X\in\mathcal{X},a\in\{0,1\}$.
Then, estimate the mediation functional by
$\hat{\psi}^{TTS}=$ $\mathbb{E}_{n}\bigg{[}\frac{I(A=1)\hat{p}(M\mid A=0,X)}{\hat{p}(A=1\mid X)\hat{p}(M\mid A=1,X)}\{Y-\hat{\mathbb{E}}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{1-\hat{p}(A=1\mid X)}\{\hat{\mathbb{E}}[Y\mid A=1,M,X]-\hat{\eta}(1,0,X)\}+\hat{\eta}(1,0,X)\bigg{]},$
where $\mathbb{E}_{n}$ is the empirical expectation operator (treating nuisance functions as non-random), and
[…]
The authors showed that this estimator is multiply robust in the following sense.
Proposition 1.
$\hat{\psi}^{TTS}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  • $\{\hat{p}(A=1\mid X),\hat{\mathbb{E}}[Y\mid A=1,M,X]\}$,
  • $\{\hat{p}(A=1\mid X),\hat{p}(M\mid A=a,X)\}$,
  • $\{\hat{\mathbb{E}}[Y\mid A=1,M,X],\hat{p}(M\mid A=a,X)\}$.
As mentioned in the introduction, there are two concerns of complexity of the nuisance functions and potential large variance about $\hat{\psi}^{TTS}$. In the next section, we propose an alternative estimation strategy that mitigates these issues.
4 Proposed Estimation Approach
We define the following nuisance functions.
$\pi_{1}(X)=\frac{1}{p(A=0\mid X)},$ $\forall X\in\mathcal{X},$
$\pi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M,A=1\mid X)},$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
$\mu_{1}(M,X)=\mathbb{E}[Y\mid A=1,M,X],$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
$\mu_{2}(X)=\mathbb{E}[\mu_{1}(M,X)\mid A=0,X],$ $\forall X\in\mathcal{X}.$
Note that the parameter of interest can be written as
$\psi_{0}=\mathbb{E}[\mu_{2}(X)].$
We start by presenting the steps of our proposed estimation method, which contains cross-fitting (Chernozhukov et al.,, 2018) for separating the samples used for estimating the nuisance functions from that used for estimating the parameter of interest.
Estimating the nuisance functions:
We partition the samples into $L$ equal size folds $\{I_{1},\dots,I_{L}\}$. For $\ell\in\{1,\dots,L\}$, we estimate the nuisance functions $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ on data from all folds but $I_{\ell}$, which we denote as $I_{\ell}^{c}$. In order for estimating the nuisance functions, we proceed in the following two stages.
  • Stage 1:
    • We estimate $\pi_{1}$ using the following balancing estimator:
      We choose $\hat{\pi}_{1,\ell}$ such that it satisfies
      $\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}h_{1}(X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1},$ (1)
      where $\mathcal{H}_{1}\subset L^{2}(P_{0,X})$ is a function space chosen by the researcher.
    • We estimate $\mu_{1}$ by regressing $Y$ on $\{M,X,A=1\}$ to obtain $\hat{\mu}_{1,\ell}$.
  • Stage 2:
    • We estimate $\pi_{2}$ using the following balancing estimator:
      Given $\hat{\pi}_{1,\ell}$ from Stage 1, we choose $\hat{\pi}_{2,\ell}$ such that it satisfies
      $\mathbb{E}\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}h_{2}(M,X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2},$ (2)
      where $\mathcal{H}_{2}\subset L^{2}(P_{0,(M,X)})$ is a function space chosen by the researcher.
    • Given $\hat{\mu}_{1,\ell}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1,\ell}$ on $\{X,A=0\}$ to obtain $\hat{\mu}_{2,\ell}$.
Estimating the parameter of interest:
For any estimates of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
$\phi(O;\tilde{\pi}_{1}$ $,\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}):=\tilde{\mu}_{2}(X)+(1-A)\tilde{\pi}_{1}(X)\{\tilde{\mu}_{1}(M,X)-\tilde{\mu}_{2}(X)\}+A\tilde{\pi}_{2}(M,X)\{Y-\tilde{\mu}_{1}(M,X)\}.$
For all $\ell$, let $\hat{\psi}^{mr}_{\ell}$ be the estimator of $\psi_{0}$ defined as,
$\hat{\psi}^{mr}_{\ell}=\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right],$
where the empirical expectation is evaluated on $I_{\ell}$. Our final estimator of the parameter of interest will be the following.
$\hat{\psi}^{mr}=\frac{1}{L}\sum_{\ell=1}^{L}\hat{\psi}^{mr}_{\ell}.$
The use of the cross-fitting approach plays a key role in establishing consistency and asymptotic normality of our proposed estimator, shown in Theorem 2.
We propose a non-parametric estimator for Stage 1 based on minimax optimization. A naive approach to implement this would be by solving the following optimization problem.
$\tilde{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right],$
where $\Pi_{1}\subset L^{2}(P_{0,X})$ is a function space chosen by the researcher. Unfortunately, this estimator will be unstable and highly sensitive to noise. The following result suggests an alternative.
Proposition 2.
$\breve{\pi}_{1}$ obtained from the following minimax optimization satisfies Equation (1) with equality.
$\breve{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(P_{0,X})}\max_{h_{1}\in L^{2}(P_{0,X})}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
Inspired by Proposition 2, we estimate $\pi_{1}$ as
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]+R_{n}^{1}(\pi_{1},h_{1}),$ (3)
where $R_{n}^{1}(\cdot,\cdot)$ is a regularizer. We will give a specific choice for $R_{n}^{1}(\cdot,\cdot)$ in Section 4.1.
Remark 1.
Our proposed nonparametric estimator $\hat{\pi}_{1}$ in Equation (3) can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect. Therefore, this provides a nonparametric version of the their covariate balancing propensity score (CBPS) method.
We can have similar arguments regarding $\pi_{2}$ in Stage 2.
Proposition 3.
For any give $\hat{\pi}_{1}$, $\breve{\pi}_{2}$ obtained from the following minimax optimization satisfies Equation (2) with equality.
$\breve{\pi}_{2}=\arg\min_{\pi_{2}\in L^{2}(P_{0,(M,X)})}\max_{h_{2}\in L^{2}(P_{0,(M,X)})}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right].$
Therefore, in Stage 2, to estimate $\pi_{2}$, we can use the following estimator.
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]+R_{n}^{2}(\pi_{2},h_{2}),$ (4)
where $\Pi_{2}\subset L^{2}(P_{0,(M,X)})$ is a user-specified function class, and $R_{n}^{2}(\cdot,\cdot)$ is a regularizer.
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, any standard parametric or non-parametric regression method can be used.
4.1 Doubly Robust Kernel Estimator
In this section, we describe solving the minimax problems (3) and (4) when RKHSes are used as the hypothesis classes. We use the method adopted by Ghassami et al., (2022), which builds upon the non-parametric estimation approach proposed by Dikkala et al., (2020) based on adversarial learning for solving integral equations within the semi-parametric proximal causal inference framework.
Let $\Pi_{1}$, $\mathcal{H}_{1}$, $\Pi_{2}$, and $\mathcal{H}_{2}$ be RKHSes with kernels $K_{\Pi_{1}}$, $K_{\mathcal{H}_{1}}$, $K_{\Pi_{2}}$ and $K_{\mathcal{H}_{2}}$, respectively, equipped with the RKHS norms $\|\cdot\|_{\Pi_{1}}$, $\|\cdot\|_{\mathcal{H}_{1}}$, $\|\cdot\|_{\Pi_{2}}$ and $\|\cdot\|_{\mathcal{H}_{2}}$. We propose the following Tikhonov regularization-based optimizations:
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$ (5)
$\hat{\pi}_{2}$ $=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\lambda_{\mathcal{H}_{2}}\|h_{2}\|_{\mathcal{H}_{2}}^{2}+\lambda_{\Pi_{2}}\|\pi_{2}\|_{\Pi_{2}}^{2},$ (6)
where $\lambda_{\mathcal{H}_{1}}$, $\lambda_{\Pi_{1}}$, $\lambda_{\mathcal{H}_{2}}$, and $\lambda_{\Pi_{2}}$ are regularization hyperparameters that are decreasing in $n$.
For the convergence analysis of the proposed minimax estimators, we direct the reader to (Dikkala et al.,, 2020; Ghassami et al.,, 2022).
Define $K_{\Pi_{1},n}=(K_{\Pi_{1}}(X_{i},X_{j}))_{i,j=1}^{n}$, $K_{\Pi_{2},n}=(K_{\Pi_{2}}((M,X)_{i},(M,X)_{j}))_{i,j=1}^{n}$, $K_{\mathcal{H}_{1},n}=(K_{\mathcal{H}_{1}}(X_{i},X_{j}))_{i,j=1}^{n}$, and $K_{\mathcal{H}_{2},n}=(K_{\mathcal{H}_{2}}((M,X)_{i},(M,X)_{j}))_{i,j=1}^{n}$ as the empirical kernel matrices corresponding to spaces $\Pi_{1}$, $\Pi_{2}$, $\mathcal{H}_{1}$, and $\mathcal{H}_{2}$, respectively. The following result provides closed-form solutions for $\hat{\pi}_{1}$ and $\hat{\pi}_{2}$.
Proposition 4.
Equation (5) achieves its optimum at $\hat{\pi}_{1}=\sum_{i=1}^{n}\alpha_{i}K_{\Pi_{1}}(X_{i},\cdot)$, with $\alpha$ defined as,
$\alpha=-\Big{(}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big{)}^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n},$
where $\Gamma_{1}=K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$, $diag(1-A)$ is a diagonal matrix with $1-A_{i}$ as the $i$-th diagonal entry, and $e_{n}:=\begin{pmatrix}1&1&\cdots&1\end{pmatrix}^{\top}$, and $\dagger$ denotes Moore-Penrose pseudo-inverse.
Equation (6) achieves its optimum at $\hat{\pi}_{2}=\sum_{i=1}^{n}\beta_{i}K_{\Pi_{2}}((M,X)_{i},\cdot)$, with $\beta$ defined as,
$\beta=-\Big{(}K_{\Pi_{2},n}diag(A)\Gamma_{2}diag(A)K_{\Pi_{2},n}+n^{2}\lambda_{\Pi_{2}}K_{\Pi_{2},n}\Big{)}^{\dagger}K_{\Pi_{2},n}diag(A)\Gamma_{2}\big{(}(1-A)\hat{\pi}_{1}(X)\big{)}_{n},$
where $\Gamma_{2}=K_{\mathcal{H}_{2},n}(\frac{1}{n}K_{\mathcal{H}_{2},n}+\lambda_{\mathcal{H}_{2}}I_{n})^{-1}$, $diag(A)$ is a diagonal matrix with $A_{i}$ as the $i$-th diagonal entry, and $\big{(}(1-A)\hat{\pi}_{1}(X)\big{)}_{n}:=\begin{pmatrix}(1-A_{1})\hat{\pi}_{1}(X_{1})&(1-A_{2})\hat{\pi}_{1}(X_{2})&\cdots&(1-A_{n})\hat{\pi}_{1}(X_{n})\end{pmatrix}^{\top}$.
5 Theoretical Analysis
We start by making the observation that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions (1) and (2) with equality.
Proposition 5.
  1. (a)
    For the true nuisance function $\pi_{1}$, we have
    $\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in L^{2}(P_{0,X}).$
  2. (b)
    For the true nuisance functions $\pi_{1}$ and $\pi_{2}$, we have
    $\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]=0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in L^{2}(P_{0,(M,X)}).$
Suppose conditions (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(P_{0,X})$ and $\mathcal{H}_{2}=L^{2}(P_{0,(M,X)})$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, when $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are not the corresponding $L^{2}$ spaces, we can have conditions (1) and (2) hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{mr}$.
Theorem 1.
The bias of the estimator $\hat{\psi}^{mr}$ satisfies the following equality:
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$ $=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
From the structure of the bias of the estimator $\hat{\psi}^{mr}$, it is evident that if there exists $h_{1}\in\mathcal{H}_{1}$ which well-approximates $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$, and there exists $h_{2}\in\mathcal{H}_{2}$ which well-approximates $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$, and further conditions (1) and (2) hold, then we can eliminate the bias of the estimator $\hat{\psi}^{mr}$ even if we do not have a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
Importantly, under the following positivity assumption, our approach retains the multiple-robustness property, as stated in Proposition 6.
Assumption 2.
For some $\epsilon>0$, we have $p(A=0\mid X)>\epsilon$ and $p(A=1\mid M,X)>\epsilon$.
Proposition 6.
Under Assumption 2 , $\hat{\psi}^{mr}$ is asymptotically unbiased if at least one of the following conditions holds:.
  • $\|\hat{\pi}_{1}(X)-\pi_{1}(X)\|_{2}=o_{p}(1),$ and $\|\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\|_{2}=o_{p}(1)$;
  • $\|\hat{\mu}_{1}(M,X)-\mu_{1}(M,X)\|_{2}=o_{p}(1),$ and $\|\hat{\mu}_{2}(X)-\mu_{2}(X)\|_{2}=o_{p}(1)$;
  • $\|\hat{\pi}_{1}(X)-\pi_{1}(X)\|_{2}=o_{p}(1),$ and $\|\hat{\mu}_{1}(M,X)-\mu_{1}(M,X)\|_{2}=o_{p}(1)$.
Figure 2 summarizes the result in Proposition 6. Note that our estimator $\hat{\psi}^{mr}$ is based on an influence function, which always leads to a second-order bias.
$\hat{\pi}_{1}$$\hat{\pi}_{2}$$\hat{\mu}_{1}$$\hat{\mu}_{2}$tage 1:Stage 2:
Figure 2: Nuisance functions estimated in each stage of the approach. The arrow from $\hat{\pi}_{1}$ to $\hat{\pi}_{2}$ indicates that the estimator of $\pi_{1}$ is used in estimating $\pi_{2}$. The arrow from $\hat{\mu}_{1}$ to $\hat{\mu}_{2}$ is interpreted similarly. Correct specification of any of the pairs of functions in an ellipse suffices unbiasedness of the final estimator of the parameter of interest.
5.1 Asymptotic Normality
We next provide sufficient conditions for consistency and asymptotic normality of the estimator $\hat{\psi}^{mr}$.
Assumption 3.
  1. (a)
    $\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}=o_{p}(1)$, $\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2}=o_{p}(1)$, $\|\mu_{1}^{2}(M,X)-\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}=o_{p}(1)$, and $\|\mu_{2}^{2}(X)-\hat{\mu}_{2,\ell}^{2}(X)\|_{2}=o_{p}(1)$.
  2. (b)
    $\|(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\|_{2}=o_{p}(1)$ and $\|(\pi_{2}(M,X)-\hat{\pi}_{2,\ell}(M,X))^{2}\|_{2}=o_{p}(1)$.
  3. (c)
    There exists some constant $C$ such that $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $\mathbb{E}[Y^{4}]$, are upper bounded by $C$.
  4. (d)
    $\big{\|}p(A=0\mid X)\hat{\pi}_{1,\ell}(X)-1\big{\|}_{2}\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2}=o_{p}(n^{-1/2})$, and
    $\big{\|}p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)\big{\|}_{2}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}=o_{p}(n^{-1/2})$.
For any choices of the nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, define,
$\Phi(O;\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2},\psi):=\phi(O;\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2})-\psi.$
We have the following result regarding the asymptotic behavior of $\hat{\psi}^{mr}$.
Theorem 2.
Under Assumptions 2 and  3, the estimator $\hat{\psi}^{mr}$ is $\sqrt{n}$-consistent and asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})=\sqrt{n}\mathbb{E}_{n}\left[\Phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2},\psi_{0})\right]+o_{p}(1).$
As a corollary of Theorem 2, one can use the influence function of $\psi_{0}$ to obtain confidence intervals for the parameter of interest. For $\ell\in\{1,\dots,L\}$, we estimate the variance of the influence function on fold $I_{\ell}$ as,
$\hat{\sigma}_{\ell}^{2}=\frac{1}{m}\sum_{i=1}^{m}\Phi^{2}(O_{i,\ell};\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell},\hat{\psi}_{\ell}^{mr}),$
where $O_{i,\ell}\in I_{\ell}$, and define,
$\hat{\sigma}^{2}=\frac{1}{L}\sum_{\ell=1}^{L}\hat{\sigma}_{\ell}^{2}.$
Then the $(1-\alpha)$ confidence interval of $\psi_{0}$ can be obtained as,
$\hat{\psi}^{mr}\pm z_{1-\alpha/2}\frac{\hat{\sigma}}{\sqrt{n}},$
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution.
6 Conclusion
In estimating a functional of interest that involves nuisance functions, if one naively aims to perfectly fit the nuisance functions, the resulting estimator of the parameter of interest can be sensitive to even small errors in the nuisance function estimators. Therefore, in order to estimate a functional, it is important to investigate the role of nuisance functions involved in that functional in the structure of the bias. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional when the influence function-based estimator is used. Our proposed estimators are based on the approach of balancing that can target the bias of the parameter of interest more directly and also can lead to an estimator which is less sensitive to error. We demonstrated that this approach retains the multiple-robustness property. Moreover, we provided conditions on our nuisance function estimators sufficient for consistency and asymptotic normality of the influence function-based estimator of the mediation functional.
Appendix A Proofs
Proof of Proposition 2.
We note that
$\arg\max_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left(\frac{1}{2}h_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}.$
Therefore,
$\breve{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[2\left\{(1-A)\pi_{1}(X)-1\right\}\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}-\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}^{2}\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}^{2}\right],$
which since zero is achievable by $\frac{1}{p(A=0\mid X)}$, it implies that
$\mathbb{E}\bigg{[}(1-A)\breve{\pi}_{1}(X)-1\bigg{|}X\bigg{]}=0.$
Therefore, for all $h_{1}\in\mathcal{H}_{1}$,
$\mathbb{E}\left[\left\{(1-A)\breve{\pi}_{1}(X)-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\mathbb{E}\bigg{[}(1-A)\breve{\pi}_{1}(X)-1\bigg{|}X\bigg{]}h_{1}(X)\right]=0.$
Proof of Proposition 3.
For any give $\hat{\pi}_{1}$, we note that
$\arg\max_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)(1-A)\hat{\pi}_{1}(X)\right\}^{2}+\frac{1}{4}h^{2}_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)\right]$
$=\arg\min_{h_{2}}\mathbb{E}\left[\left(\frac{1}{2}h_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}.$
Therefore,
$\breve{\pi}_{2}$ $=\arg\min_{\pi_{2}}\max_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{\pi_{2}}\mathbb{E}\bigg{[}2\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\mathbb{E}\big{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\big{|}M,X\big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}\big{[}A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\big{|}M,X\big{]}^{2}\bigg{]}$
$=\arg\min_{\pi_{2}}\mathbb{E}\left[\mathbb{E}\bigg{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}^{2}\right],$
which since zero is achievable by $\frac{p(M,A=0\mid X)\hat{\pi}_{1}(X)}{p(M,A=1\mid X)}$, it implies that
$\mathbb{E}\bigg{[}A\breve{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}=0.$
Therefore, for all $h_{2}\in\mathcal{H}_{2}$,
$\mathbb{E}\left[\left\{A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
$=\mathbb{E}\left[\mathbb{E}\bigg{[}A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\bigg{|}M,X\bigg{]}h_{2}(M,X)\right]=0.$
Proof of Proposition 4.
We first show that, for any function $\pi_{1}$, we have
$\max_{h_{1}\in\mathcal{H}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
=$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
where $\xi_{n}(\pi_{1})=\frac{1}{n}\big{(}(1-A_{i})\pi_{1}(X_{i})-1\big{)}_{i=1}^{n}$.
To see this, by representer theorem, the solution to the maximization will be of the form,
$h_{1}(x)=\sum_{j=1}^{n}\alpha^{\ast}_{j}K_{\mathcal{H}_{1}}(X_{j},x),$
and it suffices to find the coefficients $\alpha^{\ast}=(\alpha_{j}^{\ast})_{j=1}^{n}$.
Note that we have,
$\mathbb{E}_{n}[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)]$ $=\sum_{i=1}^{n}h(X_{i})\{\xi_{n}(\pi_{1})\}_{i}={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\},$
$\mathbb{E}_{n}[h_{1}^{2}(X)]$ $=\frac{1}{n}{\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}^{2}{\alpha^{\ast}},$
$\|h_{1}\|_{\mathcal{H}_{1}}^{2}$ $={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}{\alpha^{\ast}}.$
Therefore we have,
$\mathbb{E}_{n}\Big{[}\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big{]}-\frac{1}{4}\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\}-{\alpha^{\ast}}^{\top}(\frac{1}{4n}K_{\mathcal{H}_{1},n}^{2}+\frac{1}{4}\lambda_{\mathcal{H}_{1}}K_{\mathcal{H}_{1},n}){\alpha^{\ast}}.$
Taking the derivative and setting it to zero, the optimal coefficients are obtained as,
$\alpha^{\ast}=2(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
and we have,
$\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\Big{[}\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big{]}-\frac{1}{4}\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
=$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\}.$
Therefore, the outer minimization problem is reduced to:
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\{\xi_{n}(\pi_{1})\}^{\top}\Gamma_{1}\{\xi_{n}(\pi_{1})\}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$
where $\Gamma_{1}=K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$
By representer theorem, the solution to the maximization will be of the form,
$\pi_{1}(x)=\sum_{j=1}^{n}\alpha_{j}K_{\Pi_{1}}(X_{j},x),$
and it suffices to find the coefficients $\alpha=(\alpha_{j})_{j=1}^{n}$.
Note that,
$\xi_{n}(\pi_{1})$ $=\frac{1}{n}(diag(1-A)K_{\Pi_{1},n}\alpha-e_{n}),$
$\|\pi_{1}\|_{\Pi_{1}}^{2}$ $=\alpha^{\top}K_{\Pi_{1},n}\alpha.$
Therefore, we have,
$\min_{\pi_{1}\in\Pi_{1}}\{\xi_{n}(\pi_{1})\}^{\top}\Gamma_{1}\{\xi_{n}(\pi_{1})\}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2}$
$=\min_{\alpha\in\mathbb{R}^{n}}\frac{1}{n^{2}}\alpha^{\top}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\frac{2}{n^{2}}e_{n}^{\top}\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\lambda_{\Pi_{1}}\alpha^{\top}K_{\Pi_{1},n}\alpha+c,$
which is solved by
$\alpha=-\Big{(}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big{)}^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n}.$
The derivation for $\beta$ is similar and omitted for brevity.
Proof of Proposition 5.
  1. (a)
    For all $h_{1}\in L^{2}(P_{0,X})$, we have
    $\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\left\{\mathbb{E}[1-A\mid X]\pi_{1}(X)-1\right\}h_{1}(X)\right]=0.$
  2. (b)
    For all $h_{2}\in L^{2}(P_{0,(M,X)})$, we have
    $\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(A=1\mid M,X)p(M\mid A=0,X)}{p(M,A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(M\mid A=0,X)}{p(M\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=\mathbb{E}\left[\left\{\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}-\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=0.$
Proof of Proposition 6.
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$
$=\mathbb{E}[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}]-\psi_{0}$
$=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{\mathbb{E}\left[1-A\mid X\right]\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{\mathbb{E}\left[A\mid M,X\right]\hat{\pi}_{2}(M,X)-\mathbb{E}\left[1-A\mid M,X\right]\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{p\left(A=1\mid M,X\right)\hat{\pi}_{2}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
Using Cauchy-Schwarz inequality, we have
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$
$\leq\Big{\|}p\left(A=0\mid X\right)\hat{\pi}_{1}(X)-1\Big{\|}_{2}\Big{\|}\mu_{2}(X)-\hat{\mu}_{2}(X)\Big{\|}_{2}$
$~{}~{}~{}~{}~{}+\Big{\|}p\left(A=1\mid M,X\right)\hat{\pi}_{2}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1}(X)\Big{\|}_{2}\Big{\|}\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big{\|}_{2}$
$\leq\Big{\|}p\left(A=0\mid X\right)\Big{(}\hat{\pi}_{1}(X)-\pi_{1}(X)\Big{)}\Big{\|}_{2}\Big{\|}\mu_{2}(X)-\hat{\mu}_{2}(X)\Big{\|}_{2}$
$~{}~{}~{}~{}~{}+\Big{\|}\frac{p\left(A=1\mid M,X\right)}{\pi_{1}(X)}\Big{(}\pi_{1}(X)\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\hat{\pi}_{1}(X)\Big{)}\Big{\|}_{2}\Big{\|}\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big{\|}_{2}$
$\leq\Big{\|}\hat{\pi}_{1}(X)-\pi_{1}(X)\Big{\|}_{2}\Big{\|}\mu_{2}(X)-\hat{\mu}_{2}(X)\Big{\|}_{2}$
$~{}~{}~{}~{}~{}+\Big{(}\Big{\|}\pi_{1}(X)\Big{(}\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\Big{)}\Big{\|}_{2}+\Big{\|}\pi_{2}(M,X)\Big{(}\hat{\pi}_{1}(X)-\pi_{1}(X)\Big{)}\Big{\|}_{2}\Big{)}\Big{\|}\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big{\|}_{2}$
$\leq\Big{\|}\hat{\pi}_{1}(X)-\pi_{1}(X)\Big{\|}_{2}\Big{\|}\mu_{2}(X)-\hat{\mu}_{2}(X)\Big{\|}_{2}$
$~{}~{}~{}~{}~{}+\frac{1}{\epsilon^{2}}\Big{(}\Big{\|}\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\Big{\|}_{2}+\Big{\|}\hat{\pi}_{1}(X)-\pi_{1}(X)\Big{\|}_{2}\Big{)}\Big{\|}\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big{\|}_{2},$
where the last equation holds by Assumption 2. Therefore if at least one of the conditions in the proposition holds, we have
$\mathbb{E}[\hat{\psi}^{mr}]-\psi_{0}$ $=o_{p}(1).$
Proof of Theorem 1.
$\mathbb{E}[\hat{\theta}^{mr}]-\theta_{0}$
$=\mathbb{E}\left[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\right]-\mathbb{E}[\mu_{2}(X)]$
$=\mathbb{E}\left[\hat{\mu}_{2}(X)-\mu_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\bigg{|}X\right]-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\mathbb{E}\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\bigg{|}M,X\right]-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\mathbb{E}[(1-A)\mu_{1}(M,X)\mid X]\hat{\pi}_{1}(X)-\frac{\hat{\pi}_{1}(X)\hat{\mu}_{2}(X)}{\pi_{1}(X)}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\Big{[}\hat{\pi}_{2}(M,X)\big{(}\mathbb{E}[AY\mid M,X]-\mathbb{E}[A\mid M,X]\hat{\mu}_{1}(M,X)\big{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\Big{]}$
$=\mathbb{E}\left[\mathbb{E}[(1-A)\mid X]\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\hat{\pi}_{1}(X)-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\hat{\pi}_{2}(M,X)\mathbb{E}[A\mid M,X]\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
Proof of Theorem 2.
Assume the size of each fold of the data is $m$, that is, $n=mL$. Let $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ be the estimators of $(\pi_{1},\pi_{2},\mu_{1},\mu_{2})$ obtained on subsamples $I_{\ell}^{c}$. Then $\hat{\psi}^{mr}_{\ell}$ can be written as
$\hat{\psi}^{mr}_{\ell}=\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right].$
Therefore, we have
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})$ $=\sqrt{n}\Big{\{}\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}$
$=\frac{\sqrt{n}}{L\sqrt{m}}\sum_{\ell=1}^{L}\Bigg{\{}\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\}$
$\quad\quad\quad\quad\quad\quad\quad+\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$
$\quad\quad\quad\quad\quad\quad\quad+\sqrt{m}\Big{\{}\mathbb{E}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}\Bigg{\}},$
where for each function $f(O)$,
$\mathbb{G}_{n}[f]=\sqrt{n}\{\mathbb{E}_{n}[f]-\mathbb{E}[f]\},$
represents the empirical process. Recalling that $n=mL$, we have
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\}$ $(T_{1})$
$\quad+\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$ $(T_{2})$
$\quad+\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big{\{}\mathbb{E}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}.$ $(T_{3})$
We will show that $(T_{2})$ provides us with the term $\sqrt{n}\mathbb{E}_{n}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$, and under the assumptions of the theorem, $(T_{1})$ and $(T_{3})$ are $o_{p}(1)$.
Analysis of $(T_{1})$ Define
$A_{m}^{\ell}:=\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\},$
where the empirical processes are evaluated on $I_{\ell}$. We note that
$var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)$
$=m\cdot var\left(\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{E}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\mid I_{\ell}^{c}\right)$
$=var\big{(}\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})-\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\mid I_{\ell}^{c}\big{)}$
$\leq\mathbb{E}\big{[}\{\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})-\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\}^{2}\mid I_{\ell}^{c}\big{]}$
$=\mathbb{E}\Big{[}\Big{\{}\hat{\mu}_{2,\ell}(X)+(1-A)\hat{\pi}_{1,\ell}(X)\{\hat{\mu}_{1,\ell}(M,X)-\hat{\mu}_{2,\ell}(X)\}+A\hat{\pi}_{2,\ell}(M,X)\{Y-\hat{\mu}_{1,\ell}(M,X)\}$
$\quad\quad-\mu_{2}(X)-(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}-A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big{\}}^{2}\Big{|}I_{\ell}^{c}\Big{]}$
$=\mathbb{E}\Big{[}\Big{\{}\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
$\quad\quad-\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)-\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)-AY\pi_{2}(M,X)\Big{\}}^{2}\Big{|}I_{\ell}^{c}\Big{]}$
$\leq 5\Big{(}\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|(1-A)(\hat{\mu}_{2,\ell}(X)\hat{\pi}_{1,\ell}(X)-\mu_{2}(X)\pi_{1}(X))\|_{2}^{2}$
$\quad\quad+\|(1-A)(\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{1,\ell}(X)-\mu_{1}(M,X)\pi_{1}(X))\|_{2}^{2}$
$\quad\quad+\|A(\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{2,\ell}(M,X)-\mu_{1}(M,X)\pi_{2}(M,X))\|_{2}^{2}+\|AY(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}\Big{)}$
$\leq 5\Big{(}\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|\hat{\mu}_{2,\ell}(X)\hat{\pi}_{1,\ell}(X)-\mu_{2}(X)\pi_{1}(X)\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{1,\ell}(X)-\mu_{1}(M,X)\pi_{1}(X)\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{2,\ell}(M,X)-\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}^{2}+\|Y(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}\Big{)}$
$\leq 10\Big{(}\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|\hat{\mu}_{2,\ell}(X)(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))\|_{2}^{2}+\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))\|_{2}^{2}+\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}+\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|Y(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}\Big{)}$
$\leq 10\Big{(}\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|\hat{\mu}_{2,\ell}^{2}(X)\|_{2}\|(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))^{2}\|_{2}+\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}\|(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))^{2}\|_{2}+\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\|_{2}+\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|Y^{2}\|_{2}\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\|_{2}\Big{)}.$
By Assumption 3 (a) and (b), $\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}$, $\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}$, $\|\hat{\mu}_{1,\ell}^{2}(M,X)-\mu_{1}^{2}(M,X)\|_{2}$, $\|\hat{\mu}_{2,\ell}^{2}(X)-\mu_{2}^{2}(X)\|_{2}$, $\|(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))^{2}\|_{2}$, and $\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\|_{2}$ converge to zero in probability. By Assumption 2,
$\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}$ $\leq\frac{1}{\epsilon}\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2},$
$\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$ $\leq\frac{1}{\epsilon}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2},$
$\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$ $=\|\frac{p(A=0\mid M,X)}{p(A=1\mid M,X)}\pi_{1}(X)(\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X))\|_{2}$
$\leq\frac{1}{\epsilon^{2}}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}.$
Therefore, $\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}$, $\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$, and $\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$ also converge to zero in probability.
By Assumption 3 (c), $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $\|Y^{2}\|_{2}$ are bounded by some constants, which imply that
$\|\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}$ $\leq\|\hat{\mu}_{1,\ell}^{2}(M,X)-\mu_{1}^{2}(M,X)\|_{2}+\|\mu_{1}^{2}(M,X)\|_{2}\leq o_{p}(1)+C=O_{p}(1),$ (7)
$\|\hat{\mu}_{2,\ell}^{2}(X)\|_{2}$ $\leq\|\hat{\mu}_{2,\ell}^{2}(X)-\mu_{2}^{2}(X)\|_{2}+\|\mu_{2}^{2}(X)\|_{2}\leq o_{p}(1)+C=O_{p}(1).$ (8)
Therefore,
$var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)=o_{p}(1).$
Also note that,
$\mathbb{E}\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]$
$=\mathbb{E}\Big{[}\hat{\mu}_{2,\ell}(X)+(1-A)\hat{\pi}_{1,\ell}(X)\{\hat{\mu}_{1,\ell}(M,X)-\hat{\mu}_{2,\ell}(X)\}+A\hat{\pi}_{2,\ell}(M,X)\{Y-\hat{\mu}_{1,\ell}(M,X)\}$
$\quad\quad-\mu_{2}(X)-(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}-A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big{|}I_{\ell}^{c}\Big{]}$
$=\mathbb{E}\Big{[}\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
$\quad\quad-\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)-\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)-AY\pi_{2}(M,X)\Big{|}I_{\ell}^{c}\Big{]}$
$=\mathbb{E}\Big{[}\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)\left(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)\right)$
$\quad\quad-\mu_{2}(X)\underbrace{\left(1-p(A=0\mid X)\pi_{1}(X)\right)}_{=0}$
$\quad\quad-\mu_{1}(M,X)\underbrace{\left(p(A=0\mid M,X)\pi_{1}(X)-p(A=1\mid M,X)\pi_{2}(M,X)\right)}_{=0}$
$\quad\quad+\mathbb{E}[AY\mid M,X]\hat{\pi}_{2,\ell}(M,X)-\mathbb{E}[AY\mid M,X]\pi_{2}(M,X)\Big{|}I_{\ell}^{c}\Big{]}$
$=\mathbb{E}\Big{[}\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)\left(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)\right)$
$\quad\quad+p(A=1\mid M,X)\mu_{1}(M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=1\mid M,X)\mu_{1}(M,X)\pi_{2}(M,X)\Big{|}I_{\ell}^{c}\Big{]}.$
By Jensen’s inequality and Cauchy-Schwarz inequality,
$\mathbb{E}\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]$
$\leq\Big{\|}\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)\left(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)\right)$
$\quad\quad+p(A=1\mid M,X)\mu_{1}(M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=1\mid M,X)\mu_{1}(M,X)\pi_{2}(M,X)\Big{\|}_{2}$
$=\Big{\|}\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=0\mid M,X)\pi_{1}(X)$
$\quad\quad\quad+p(A=1\mid M,X)\pi_{2}(M,X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X))$
$\quad\quad+p(A=1\mid M,X)\mu_{1}(M,X)(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\Big{\|}_{2}$
$\leq\Big{\|}\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)\Big{\|}_{2}$
$\quad\quad+\Big{\|}\hat{\mu}_{1,\ell}(M,X)p(A=0\mid M,X)\left(\hat{\pi}_{1,\ell}(X)-\pi_{1,\ell}(X)\right)\Big{\|}_{2}$
$\quad\quad+\Big{\|}\mu_{1,\ell}(M,X)p(A=1\mid M,X)(\pi_{2}(M,X)-\hat{\pi}_{2,\ell}(M,X))\Big{\|}_{2}$
$\quad\quad+\Big{\|}p(A=1\mid M,X)\mu_{1}(M,X)(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\Big{\|}_{2}$
$\leq\Big{\|}\hat{\mu}_{2,\ell}^{2}(X)\Big{\|}_{2}^{1/2}\Big{\|}(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}(p(A=0\mid M,X)\hat{\mu}_{1,\ell}(M,X))^{2}\Big{\|}_{2}^{1/2}\Big{\|}\big{(}\hat{\pi}_{1,\ell}(X)-\pi_{1}(X)\big{)}^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}(p(A=1\mid M,X)\hat{\mu}_{1,\ell}(M,X))^{2}\Big{\|}_{2}^{1/2}\Big{\|}\big{(}\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X)\big{)}^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}(p(A=1\mid M,X)\mu_{1}(M,X))^{2}\Big{\|}_{2}^{1/2}\Big{\|}(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\Big{\|}_{2}^{1/2}$
$\leq\Big{\|}\hat{\mu}_{2,\ell}^{2}(X)\Big{\|}_{2}^{1/2}\Big{\|}(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}\hat{\mu}_{1,\ell}^{2}(M,X)\Big{\|}_{2}^{1/2}\Big{\|}\big{(}\hat{\pi}_{1,\ell}(X)-\pi_{1}(X)\big{)}^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}\hat{\mu}_{1,\ell}^{2}(M,X)\Big{\|}_{2}^{1/2}\Big{\|}\big{(}\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X)\big{)}^{2}\Big{\|}_{2}^{1/2}$
$\quad\quad+\Big{\|}\mu_{1}^{2}(M,X)\Big{\|}_{2}^{1/2}\Big{\|}(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\Big{\|}_{2}^{1/2}.$
Also note that,
$\Big{\|}(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big{\|}_{2}$ $=\Big{\|}(p(A=0\mid X))^{2}(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\Big{\|}_{2}$
$\leq\Big{\|}(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\Big{\|}_{2}.$
Hence, by Assumption 3 (b), $\Big{\|}(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big{\|}_{2}$ converges to zero in probability.
Similar to inequalities (7) and (8), $\Big{\|}\hat{\mu}_{1,\ell}^{2}(M,X)\Big{\|}_{2}^{1/2}$ and $\Big{\|}\hat{\mu}_{2,\ell}^{2}(X)\Big{\|}_{2}^{1/2}$ are $O_{p}(1)$. Therefore, by Assumption 3 (a), (b) and (c),
$\mathbb{E}\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]=o_{p}(1).$
Therefore,
$\mathbb{E}\left[(A_{m}^{\ell})^{2}\mid I_{\ell}^{c}\right]=var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)+\Big{(}\mathbb{E}\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]\Big{)}^{2}\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
Then by conditional Chebyshev inequality, for all $\delta>0$, we have,
$P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
That is, $P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})$ is a bounded sequence that converges to zero in probability. Therefore,
$\mathbb{E}\left[P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})\right]=P(|A_{m}^{\ell}|>\delta)\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
That is, $A_{m}^{\ell}\stackrel{{\scriptstyle p.}}{{\rightarrow}}0$ as $m\rightarrow\infty$.
The conclusion holds for all $\ell\in\{1,\dots,L\}$ and hence, we conclude that $T_{1}=o_{p}(1)$.
Analysis of $T_{2}$
$T_{2}$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{E}_{m}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]-\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\frac{1}{m}\sum_{\ell=1}^{L}\sum_{i=1}^{m}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}.$
By central limit theorem, the last expression converges to the distribution $\mathcal{N}(0,var(\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2}))$ if $\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$. Note that $\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$ can be upper bounded as
$\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$
$=\mathbb{E}\Big{[}\Big{\{}\mu_{2}(X)+(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}+A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big{\}}^{2}\Big{]}$
$=\mathbb{E}\Big{[}\Big{\{}\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)+\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)+AY\pi_{2}(M,X)\Big{\}}^{2}\Big{]}$
$=\|\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)+\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)+AY\pi_{2}(M,X)\|_{2}^{2}$
$\leq\Big{(}\|\mu_{2}(X)\big{(}1-(1-A)\pi_{1}(X)\big{)}\|_{2}+\|\mu_{1}(M,X)\big{(}(1-A)\pi_{1}(X)-A\pi_{2}(M,X)\big{)}\|_{2}$
$\quad\quad+\|AY\pi_{2}(M,X)\|_{2}\Big{)}^{2}$
$\leq\Big{(}\|\mu_{2}(X)\|_{2}+\|(1-A)\mu_{2}(X)\pi_{1}(X)\|_{2}+\|(1-A)\mu_{1}(M,X)\pi_{1}(X)\|_{2}+\|A\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}$
$\quad\quad+\|AY\pi_{2}(M,X)\|_{2}\Big{)}^{2}$
$\leq\Big{(}\|\mu_{2}(X)\|_{2}+\|\mu_{2}(X)\pi_{1}(X)\|_{2}+\|\mu_{1}(M,X)\pi_{1}(X)\|_{2}+\|\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}+\|Y\pi_{2}(M,X)\|_{2}\Big{)}^{2}$
By Assumption 2 and Jensen’s inequality,
$\|\mu_{2}(X)\|_{2}\leq\|\mu_{2}^{2}(X)\|_{2}^{1/2},$
$\|\mu_{2}(X)\pi_{1}(X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{2}(X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{2}^{2}(X)\|_{2}^{1/2},$
$\|\mu_{1}(M,X)\pi_{1}(X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{1}(M,X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{1}^{2}(M,X)\|_{2}^{1/2},$
$\|\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}\leq\frac{1}{\epsilon^{2}}\|\mu_{1}(M,X)\|_{2}\leq\frac{1}{\epsilon^{2}}\|\mu_{1}^{2}(M,X)\|_{2}^{1/2},$
$\|Y\pi_{2}(M,X)\|_{2}\leq\frac{1}{\epsilon^{2}}\|Y\|_{2}\leq\frac{1}{\epsilon^{2}}\|Y^{2}\|_{2}^{1/2}.$
Therefore, by Assumption 3 (c), we have $\mathbb{E}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$.
Analysis of $T_{3}$
Using result of Theorem 1 and Law of Total Probability, we have,
$T_{3}$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big{\{}\mathbb{E}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{E}\left[\hat{\psi}^{mr}_{\ell}\right]-\psi_{0}\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~{}~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\mathbb{E}\left[\left\{\mathbb{E}\left[1-A\mid X\right]\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~{}~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{\mathbb{E}\left[A\mid M,X\right]\hat{\pi}_{2,\ell}(M,X)-\mathbb{E}\left[1-A\mid M,X\right]\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\mathbb{E}\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathbb{E}[\{p\left(A=1\mid M,X\right)\hat{\pi}_{2,\ell}(M,X)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-p\left(A=0\mid M,X\right)\hat{\pi}_{1,\ell}(X)\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}]\Bigg{\}}.$
Using Cauchy-Schwarz inequality, we have
$T_{3}$ $\leq\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\Big{\|}p\left(A=0\mid X\right)\hat{\pi}_{1,\ell}(X)-1\Big{\|}_{2}\Big{\|}\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\Big{\|}_{2}$
$~{}~{}~{}~{}~{}~{}+\Big{\|}p\left(A=1\mid M,X\right)\hat{\pi}_{2,\ell}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1,\ell}(X)\Big{\|}_{2}\Big{\|}\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\Big{\|}_{2}\Bigg{\}}.$
Therefore by Assumption 3 (d), we have
$T_{3}$ $=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}o_{p}(n^{-1/2})+o_{p}(n^{-1/2})\Bigg{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\times L\times o_{p}(n^{-1/2})$
$=\sqrt{mL}\times o_{p}(n^{-1/2})$
$=\sqrt{n}o_{p}(n^{-1/2})$
$=o_{p}(1).$