### v3

# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

Abstract
Tchetgen Tchetgen and Shpitser (2012) introduced an efficient, debiased, and robust influence function-based estimator for the mediation functional, which is the key component in mediation analysis. This estimator relies on the treatment, mediator, and outcome mean mechanisms. However, treating these three mechanisms as nuisance functions and fitting them as accurately as possible may not be the most effective approach. Instead, it is essential to identify the specific functionals and aspects of these mechanisms that impact the estimation of the mediation functional. In this work, we consider a specific reparametrization of the likelihood function that requires four nuisance functions. To estimate them, we propose a two-stage estimation strategy guided by the role of the nuisance functions in the bias structure of the influence function-based estimator. In particular, two of the functions are estimated using a novel nonparametric weighted balancing approach that directly targets the bias of the final mediation functional estimator. We show that the resulting estimator is consistent and asymptotically normal under certain conditions and attains multiple robustness against misspecifications of the nuisance functions. In simulations, our estimator demonstrated better stability and reduced bias and mean squared error compared to the original influence function-based estimator and a naive estimator. In an application to NHANES 2013-2014 data, our approach suggests that obesity results in a 60% higher odds of coronary heart disease, with roughly 35% higher odds attributable to the direct pathway independent of Glycohemoglobin.
Keywords: Balancing Estimators; Causal Mediation Analysis; Influence Function-Based Estimation; Multiple Robustness.
1 Introduction
In contrast to merely assessing the overall effect of a treatment variable on an outcome variable of interest, causal mediation analysis delves deeper by investigating the underlying mechanisms through which the treatment influences the outcome. This unique focus has made it an indispensable tool across diverse fields, including social science, economics, and epidemiology (VanderWeele,, 2015; MacKinnon,, 2012; Hayes,, 2017; Richiardi et al.,, 2013; Ten Have and Joffe,, 2012). The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model, the direct and indirect effects can be identified using a functional of the observed distribution known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those conditions are met in the setting under study and the parameter is identified, the next question is how to estimate the mediation functional.
There has been several estimation strategies proposed in the literature based on regression and weighting techniques (Imai et al.,, 2010; Lange et al.,, 2012; Valeri and VanderWeele,, 2013; VanderWeele et al.,, 2014; Díaz and Hejazi,, 2020; Hejazi et al.,, 2023; Rudolph et al.,, 2023). In their seminal work, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric efficient and has the desired property that, under certain regularity conditions, remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism (i.e., the conditional distribution of the treatment given the covariates), one describing the mediator mechanism (i.e., the conditional distribution of the mediator variable given covariates and the treatment), and one describing the outcome mean mechanism (i.e., the conditional expectation of the outcome variable given covariates, the treatment, and the mediator), and the model for two out of three should be correctly specified. There are three challenges with their proposed estimator: (1) among the aforementioned nuisance functions, the mediator mechanism is usually a difficult object to estimate, as it may be the density of a multivariate continuous variable conditioned on another multivariate continuous variable; (2) in their estimator of the mediation functional, the estimators of the treatment and mediator mechanisms appear in the denominator, and hence can significantly inflate the variance of the estimator, especially when these quantities are small; (3) their choice of nuisance functions is not explicitly guided by the role of nuisance functions in the bias structure of the estimator of the causal parameter of interest. In a later work, Díaz et al., (2021) considered a reparametrization of the likelihood function to avoid the estimation of the mediation mechanism. [^1] Their framework additionally allows for mediator-outcome confounders affected by treatment. Instead, it relies on an additional nuisance function describing the conditional distribution of the treatment given the covariates and the mediator. While their approach also achieves multiple robustness, it shares the same drawbacks as the original formulation of Tchetgen Tchetgen and Shpitser, (2012)’s estimator: their new nuisance function is difficult to estimate in high-dimensional settings, will appear in denominators leading to large variance, and may even be inconsistent with the treatment mechanism under parametric modeling. Moreover, their approach also fails to consider the bias structure of the estimator. In both methods of Tchetgen Tchetgen and Shpitser, (2012) and Díaz et al., (2021), trying to naively fit the nuisance functions as good as possible will not necessarily lead to an estimator with reasonable performance. Instead, it is important to understand what aspects of the nuisance functions should be focused on to obtain a better estimate of the mediation functional.
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis. Our method can be viewed as a reparametrization of the likelihood function, designed to address the estimation challenges discussed above. The key advantage of this reparametrization is that it avoids direct estimation of the treatment and mediator mechanisms. In the first stage, we estimate two basic nuisance functions, namely the inverse of the treatment mechanism and the outcome mean mechanism. In the second stage, these basic nuisance functions are used to construct two additional nuisance functions that contain the needed information regarding the mediator mechanism. We propose a nonparametric weighted balancing approach to design the estimator of the inverse of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide an implementation of our nonparametric estimators with reproducing kernel Hilbert spaces (RKHS) as the function classes, and provide closed-form solutions for the nuisance functions. The other two nuisance functions are obtained using standard nonparametric regressions. After obtaining estimates of all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain an asymptotically unbiased estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest. We evaluate our proposed approach on both synthetic and real data. In Section 7, we apply our method to estimate the effect of obesity on coronary heart disease mediated by Glycohemoglobin, using data from the 2013-2014 cycle of National Health and Nutrition Examination Survey (National Center for Health Statistics,, 2014).
2 Model Description
AAXXMMYY
Figure 1: A graphical causal model for the mediation model.
In this section, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. We consider a model with binary treatment variable $A\in\{0,1\}$, pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $p(m\mid a,x)>0$, and $p(a\mid x)>0$.
We also require that the model does not contain any unobserved confounders, which is a common and standard assumption in the causal mediation analysis literature. This can be formalized using the sequential exchangeability assumption (Imai et al.,, 2010).
Assumption 1 (Sequential exchangeability).
For any $a,a^{\prime}\in\{0,1\}$, and any value of the mediator $m$, we have
  1. (i)
    $Y^{(a,m)}\perp\mkern-9.5mu\perp\{A,M\}\mid X$,
  2. (ii)
    $M^{(a)}\perp\mkern-9.5mu\perp A\mid X$,
  3. (iii)
    $Y^{(a,m)}\perp\mkern-9.5mu\perp M^{(a^{\prime})}\mid X$.
Intuitively, Part (i) of the assumption requires that the treatment-outcome and mediator-outcome relations are not confounded by unobserved variables, Part (ii) requires that the treatment-mediator relation is not confounded by unobserved variables (see Figure 1), and Part (iii), which is a so-called cross-world independence assumption, requires that the potential outcomes of $Y$ and $M$ even for different assignments of the treatment are conditionally independent. All three parts are untestable and should be decided based on expert’s knowledge and the researcher’s belief about how the variables in the system are generated.
Consider the parameter average treatment effect (ATE) of the treatment $A$ on the outcome $Y$, represented by $\theta_{ATE}=E[Y^{(1)}-Y^{(0)}]$. The main goal in causal mediation analysis is to quantify the part of the ATE that is mediated by the variable $M$, and the part that is directly reaching the outcome variable. To this end, the ATE can be partitioned as follows (Robins and Greenland,, 1992; Pearl,, 2001),
$\theta_{ATE}$ $=E[Y^{(1)}-Y^{(0)}]$
$=E[Y^{(1,M^{(1)})}-Y^{(0,M^{(0)})}]$
$=\underbrace{E[Y^{(1,M^{(1)})}-Y^{(1,M^{(0)})}]}_{\text{NIE}}+\underbrace{E[Y^{(1,M^{(0)})}-Y^{(0,M^{(0)})}]}_{\text{NDE}}.$
The first and the second terms in the last expression are called the natural indirect effect (NIE) and the natural direct effect (NDE) of the treatment on the outcome, respectively (Pearl,, 2001). [^2] These two terms are called the total indirect effect and the pure direct effect, respectively by Robins and Greenland, (1992). NIE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is fixed at $a=1$, while the mediator behaves as if the treatment had been changed from value $a=0$ to $a=1$. NDE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is changed from value $a=0$ to $a=1$, while the mediator behaves as if the treatment is fixed at $a=0$.
We notice that under consistency, positivity, and sequential exchangeability, we have
$E[Y^{(1,M^{(1)})}]=E[Y^{(1)}]=E\{E[Y^{(1)}\mid X]\}=E\{E[Y^{(1)}\mid A=1,X]\}=E\{E[Y\mid A=1,X]\}.$
That is $E[Y^{(1,M^{(1)})}]$ is identified. Similar argument holds for $E[Y^{(0,M^{(0)})}]$. The parameter $\theta_{0}=E[Y^{(1,M^{(0)})}]$ requires more caution. It is known that under the assumptions of our model, this parameter can be identified as follows (Pearl,, 2001; Imai et al.,, 2010).
$\theta_{0}$ $=E[Y^{(1,M^{(0)})}]$
$=\int yp(Y^{(1,m)}=y\mid x)p(M^{(0)}=m\mid x)p(x)dydmdx$
$=\int yp(y\mid a=1,m,x)p(m\mid a=0,x)p(x)dydmdx$
$=:\psi_{0}.$
The functional $\psi_{0}$ is called the mediation functional (also known as mediation formula). This quantity is purely a functional of the observed data distribution, $P_{0}$. Note that under the identification assumptions in our model, $\psi_{0}=\theta_{0}$, but in general, these two parameters are not necessarily equal. In the sequel, we refer to $\psi_{0}$ as the parameter of interest. Our goal is to estimate this parameter of interest from the independent and identically distributed observations $\{O_{i}:=(X,A,M,Y)_{i}\}_{i=1}^{n}$ drawn from $P_{0}$.
3 Influence Function-Based Estimator
In this section, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. The authors showed that in a nonparametric model, the efficient influence function of the mediation functional is given by
$IF_{\psi_{0}}=$ $\frac{I(A=1)p(M\mid A=0,X)}{p(A=1\mid X)p(M\mid A=1,X)}\{Y-E[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{p(A=0\mid X)}\{E[Y\mid A=1,M,X]-\eta(1,0,X)\}+\eta(1,0,X)-\psi_{0},$
where
$\eta(a^{\prime},a,X)=\int E[Y\mid A=a^{\prime},M=m,X]p(M=m\mid A=a,X)dm.$
Based on $IF_{\psi_{0}}$, they proposed to first estimate the following nuisance functions.
  1. (i)
    $p(A=1\mid X=\cdot)$,
  2. (ii)
    $E[Y\mid A=1,M=\cdot,X=\cdot]$,
  3. (iii)
    $p(M=\cdot\mid A=a,X=\cdot)$, for $a\in\{0,1\}$.
Then, estimate the mediation functional by
$\hat{\psi}^{\text{TTS}}=$ $E_{n}\bigg[\frac{I(A=1)\hat{p}(M\mid A=0,X)}{\hat{p}(A=1\mid X)\hat{p}(M\mid A=1,X)}\{Y-\hat{E}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{1-\hat{p}(A=1\mid X)}\{\hat{E}[Y\mid A=1,M,X]-\hat{\eta}(1,0,X)\}+\hat{\eta}(1,0,X)\bigg],$
where $E_{n}$ is the empirical expectation operator (treating nuisance functions as non-random), and
$\hat{\eta}(1,0,X)=\int\hat{E}[Y\mid A=1,M=m,X]\hat{p}(M=m\mid A=0,X)dm.$
The authors showed that this estimator is multiply robust in the following sense,
Proposition 1.
$\hat{\psi}^{\text{TTS}}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  1. (i)
    $\{\hat{p}(A=1\mid X=\cdot),\hat{E}[Y\mid A=1,M=\cdot,X=\cdot]\}$,
  2. (ii)
    $\{\hat{p}(A=1\mid X=\cdot),\hat{p}(M=\cdot\mid A=a,X=\cdot)\}$,
  3. (iii)
    $\{\hat{E}[Y\mid A=1,M=\cdot,X=\cdot],\hat{p}(M=\cdot\mid A=a,X=\cdot)\}$.
As mentioned in the introduction, the main concerns with $\hat{\psi}^{\text{TTS}}$ are the complexity of the nuisance functions and variance inflation from treatment and mediator mechanisms in the denominator. In the next section, we propose an alternative estimation strategy that mitigates these issues.
4 Proposed Estimation Approach
4.1 Two-Stage Nuisance Function Estimation
In this section, we present our two-stage nuisance function estimation approach. We start by introducing a reparametrization of the likelihood function based on the following nuisance functions.
$\pi_{1}(X=\cdot):=\frac{1}{p(A=0\mid X=\cdot)},$
$\pi_{2}(M=\cdot,X=\cdot):=\frac{p(M=\cdot\mid A=0,X=\cdot)}{p(M=\cdot,A=1\mid X=\cdot)},$
$\mu_{1}(M=\cdot,X=\cdot):=E[Y\mid A=1,M=\cdot,X=\cdot],$
$\mu_{2}(X=\cdot):=E[\mu_{1}(M,X)\mid A=0,X=\cdot].$
Note that the parameter of interest can be written as
$\psi_{0}=E[\mu_{2}(X)].$
We start by presenting the steps of our proposed estimation method, which contain cross-fitting (Chernozhukov et al.,, 2018) for separating the samples used for estimating the nuisance functions from that used for estimating the parameter of interest.
Estimating the nuisance functions. We partition the samples into $L$ folds $\{I_{1},\dots,I_{L}\}$ with roughly equal size $m=n/L$. For $\ell\in\{1,\dots,L\}$, we estimate the nuisance functions as $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ on data from all folds but $I_{\ell}$, which we denote as $I_{\ell}^{c}$. In order to estimate the nuisance functions, we proceed in the following two stages.
Stage 1:
  1. (i)
    We estimate $\pi_{1}$ using the following balancing estimator:
    We choose $\hat{\pi}_{1,\ell}$ such that it satisfies
    $E\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}h_{1}(X)\right]\approx 0,~~~~~~\forall h_{1}\in\mathcal{H}_{1},$ (1)
    where $\mathcal{H}_{1}\subset L^{2}(P_{0,X})$ is a function space chosen by the researcher.
  2. (ii)
    We estimate $\mu_{1}$ by regressing $Y$ on $\{M,X,A=1\}$ to obtain $\hat{\mu}_{1,\ell}$.
Stage 2:
  1. (i)
    We estimate $\pi_{2}$ using the following balancing estimator:
    Given $\hat{\pi}_{1,\ell}$ from Stage 1, we choose $\hat{\pi}_{2,\ell}$ such that it satisfies
    $E\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}h_{2}(M,X)\right]\approx 0,~~~~~~\forall h_{2}\in\mathcal{H}_{2},$ (2)
    where $\mathcal{H}_{2}\subset L^{2}(P_{0,(M,X)})$ is a function space chosen by the researcher.
  2. (ii)
    Given $\hat{\mu}_{1,\ell}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1,\ell}(X,M)$ on $\{X,A=0\}$ to obtain $\hat{\mu}_{2,\ell}$.
Estimating the parameter of interest. For any choice of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
$\phi(O;\tilde{\pi}_{1}$ $,\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}):=\tilde{\mu}_{2}(X)+(1-A)\tilde{\pi}_{1}(X)\{\tilde{\mu}_{1}(M,X)-\tilde{\mu}_{2}(X)\}+A\tilde{\pi}_{2}(M,X)\{Y-\tilde{\mu}_{1}(M,X)\}.$
For all $\ell$, let $\hat{\psi}^{\text{2S}}_{\ell}$ be the estimator of $\psi_{0}$ defined as,
$\hat{\psi}^{\text{2S}}_{\ell}=E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right],$
where the empirical expectation is evaluated on $I_{\ell}$. Our final estimator of the parameter of interest will be the following.
$\hat{\psi}^{\text{2S}}=\frac{1}{L}\sum_{\ell=1}^{L}\hat{\psi}^{\text{2S}}_{\ell}.$
The use of the cross-fitting approach plays a key role in establishing consistency and asymptotic normality of our proposed estimator, shown in Theorem 2.
We propose a nonparametric estimator for Stage 1 based on minimax optimization. To make the notation simple, throughout the rest of Section 4, assume we use a sample of size $n$ for estimating the nuisance functions. A direct implementation is to solve the following optimization problem.
$\tilde{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right],$
where $\Pi_{1}\subset L^{2}(P_{0,X})$ is a function space chosen by the researcher. Unfortunately, this estimator will be unstable and highly sensitive to noise. The following result suggests an alternative.
Proposition 2.
$\breve{\pi}_{1}$ obtained from the following minimax optimization satisfies Equation (1) with equality.
$\breve{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(P_{0,X})}\max_{h_{1}\in L^{2}(P_{0,X})}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
Inspired by Proposition 2, we estimate $\pi_{1}$ as,
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]+R_{n}^{1}(\pi_{1},h_{1}),$ (3)
where $R_{n}^{1}(\cdot,\cdot)$ is a regularizer. We will give a specific choice for $R_{n}^{1}(\cdot,\cdot)$ in Section 4.2.
Remark 1.
Imai and Ratkovic, (2014) proposed the covariate balancing propensity score (CBPS) method for estimating the average treatment effect. Our proposed technique for estimating $\hat{\pi}_{1}$ in Equation (3) can be viewed as a nonparametric version of their covariate balancing approach. Therefore, our proposed nonparametric estimator $\hat{\pi}_{1}$ can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect without resorting to parametric models for the inverse of the propensity score.
We have similar arguments regarding $\pi_{2}$ in Stage 2, which lead to the following result.
Proposition 3.
For any give $\hat{\pi}_{1}$, $\breve{\pi}_{2}$ obtained from the following minimax optimization satisfies Equation (2) with equality.
$\breve{\pi}_{2}=\arg\min_{\pi_{2}\in L^{2}(P_{0,(M,X)})}\max_{h_{2}\in L^{2}(P_{0,(M,X)})}E\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right].$
Therefore, in Stage 2, to estimate $\pi_{2}$, we can use the following estimator.
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}E_{n}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]+R_{n}^{2}(\pi_{2},h_{2}),$ (4)
where $\Pi_{2}\subset L^{2}(P_{0,(M,X)})$ is a user-specified function class, and $R_{n}^{2}(\cdot,\cdot)$ is a regularizer.
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, one may use any standard parametric or nonparametric regression method.
4.2 A Kernel Method-Based Implementation
In this section, we describe solving the minimax problems (3) and (4) when RKHSes are used as the hypothesis classes. We use the method of Ghassami et al., (2022), which extends the nonparametric estimation approach proposed by Dikkala et al., (2020) for solving integral equations via adversarial learning within the semiparametric proximal causal inference framework.
Let $\Pi_{1}$, $\mathcal{H}_{1}$, $\Pi_{2}$, and $\mathcal{H}_{2}$ be RKHSes with kernels $K_{\Pi_{1}}$, $K_{\mathcal{H}_{1}}$, $K_{\Pi_{2}}$, and $K_{\mathcal{H}_{2}}$, respectively, equipped with the RKHS norms $\|\cdot\|_{\Pi_{1}}$, $\|\cdot\|_{\mathcal{H}_{1}}$, $\|\cdot\|_{\Pi_{2}}$, and $\|\cdot\|_{\mathcal{H}_{2}}$. We propose the following Tikhonov regularization-based optimization problems:
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$ (5)
$\hat{\pi}_{2}$ $=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}E_{n}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-\lambda_{\mathcal{H}_{2}}\|h_{2}\|_{\mathcal{H}_{2}}^{2}+\lambda_{\Pi_{2}}\|\pi_{2}\|_{\Pi_{2}}^{2},$ (6)
where $\lambda_{\mathcal{H}_{1}}$, $\lambda_{\Pi_{1}}$, $\lambda_{\mathcal{H}_{2}}$, and $\lambda_{\Pi_{2}}$ are regularization hyperparameters that are decreasing in $n$. The convergence analysis of the proposed minimax estimators can be done similar to the analysis presented in (Dikkala et al.,, 2020; Ghassami et al.,, 2022), and hence omitted here.
Define $K_{\Pi_{1},n}=(K_{\Pi_{1}}(X_{i},X_{j}))_{i,j=1}^{n}$, $K_{\Pi_{2},n}=(K_{\Pi_{2}}((M,X)_{i},(M,X)_{j}))_{i,j=1}^{n}$, $K_{\mathcal{H}_{1},n}=(K_{\mathcal{H}_{1}}(X_{i},X_{j}))_{i,j=1}^{n}$, and $K_{\mathcal{H}_{2},n}=(K_{\mathcal{H}_{2}}((M,X)_{i},(M,X)_{j}))_{i,j=1}^{n}$ as the empirical kernel matrices corresponding to spaces $\Pi_{1}$, $\Pi_{2}$, $\mathcal{H}_{1}$, and $\mathcal{H}_{2}$, respectively. The following result provides closed-form solutions for $\hat{\pi}_{1}$ and $\hat{\pi}_{2}$.
Proposition 4.
  1. (a)
    Equation (5) achieves its optimum at $\hat{\pi}_{1}=\sum_{i=1}^{n}\alpha_{i}K_{\Pi_{1}}(X_{i},\cdot)$, with $\alpha$ defined as,
    $\alpha=\Big(K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big)^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n},$
    where $\Gamma_{1}=\frac{1}{4}K_{\mathcal{H}_{1},n}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$, diag(1A)diag(1-A) is a diagonal matrix with $1-A_{i}$ as the $i$-th diagonal entry, $e_{n}:=\begin{pmatrix}1&1&\cdots&1\end{pmatrix}^{\top}$, and $\dagger$ denotes Moore-Penrose pseudo-inverse.
  2. (b)
    Equation (6) achieves its optimum at $\hat{\pi}_{2}=\sum_{i=1}^{n}\beta_{i}K_{\Pi_{2}}((M,X)_{i},\cdot)$, with $\beta$ defined as,
    $\beta=\Big(K_{\Pi_{2},n}diag(A)\Gamma_{2}diag(A)K_{\Pi_{2},n}+n^{2}\lambda_{\Pi_{2}}K_{\Pi_{2},n}\Big)^{\dagger}K_{\Pi_{2},n}diag(A)\Gamma_{2}\big((1-A)\hat{\pi}_{1}(X)\big)_{n},$
    where $\Gamma_{2}=\frac{1}{4}K_{\mathcal{H}_{2},n}(\frac{1}{4n}K_{\mathcal{H}_{2},n}+\lambda_{\mathcal{H}_{2}}I_{n})^{-1}$, $diag(A)$ is a diagonal matrix with $A_{i}$ as the $i$-th diagonal entry, and $\big((1-A)\hat{\pi}_{1}(X)\big)_{n}:=\begin{pmatrix}(1-A_{1})\hat{\pi}_{1}(X_{1})&(1-A_{2})\hat{\pi}_{1}(X_{2})&\cdots&(1-A_{n})\hat{\pi}_{1}(X_{n})\end{pmatrix}^{\top}$.
5 Theoretical Analysis
5.1 Bias Analysis and Robustness Properties
We begin by showing that the true nuisance functions $\pi_{1}$ and $\pi_{2}$ satisfy conditions (1) and (2) with equality, as stated in the following result.
Proposition 5.
  1. (a)
    For the true nuisance function $\pi_{1}$, we have
    $E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=0,~~~~~~\forall h_{1}\in L^{2}(P_{0,X}).$
  2. (b)
    For the true nuisance functions $\pi_{1}$ and $\pi_{2}$, we have
    $E\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]=0,~~~~~~\forall h_{2}\in L^{2}(P_{0,(M,X)}).$
Suppose conditions (1) and (2) hold with equality and $\mathcal{H}_{1}=L^{2}(P_{0,X})$ and $\mathcal{H}_{2}=L^{2}(P_{0,(M,X)})$. Then, it is easy to see that $\hat{\pi}_{1}=\pi_{1}$ and $\hat{\pi}_{2}=\pi_{2}$ with probability one. However, if $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are restricted hypothesis spaces rather than the full $L^{2}$ spaces, conditions (1) and (2) may still hold with equality even though $\hat{\pi}_{1}\neq\pi_{1}$ and $\hat{\pi}_{2}\neq\pi_{2}$. Nevertheless, satisfying conditions (1) and (2) can be sufficient for controlling the bias of the final estimator of the parameter of interest. We formalize this argument by looking at the structure of the bias of $\hat{\psi}^{\text{2S}}$.
Theorem 1.
The bias of the estimator $\hat{\psi}^{\text{2S}}$ satisfies the following equality:
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$ $=E\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
Note that our estimator $\hat{\psi}^{\text{2S}}$ is based on an influence function, which always leads to a second-order bias. From the structure of the bias of the estimator $\hat{\psi}^{\text{2S}}$, it is evident that if there exists $h_{1}\in\mathcal{H}_{1}$ which closely approximates $\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}$, and there exists $h_{2}\in\mathcal{H}_{2}$ which closely approximates $\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}$, and further conditions (1) and (2) hold, then we can eliminate the bias of the estimator $\hat{\psi}^{\text{2S}}$ even if we do not have a perfect fit of $\pi_{1}$ and $\pi_{2}$. This is in fact the main motivation of our two-stage approach. It substitutes fitting $p(A=1\mid X)$ and $p(M\mid A,X)$, which can be a difficult task, by a much weaker requirement based on balancing.
We, next, analyze the robustness of our proposed method against misspecification of the nuisance functions. Under the following positivity assumption, our approach possesses a multiple-robustness property, as stated in Proposition 6.
Assumption 2.
For some $\epsilon>0$, we have $p(A=a\mid X=x)>\epsilon$ and $p(M=m\mid A=a,X=x)>\epsilon$, for all $m$, $a$, and $x$.
Proposition 6.
Under Assumption 2 , $\hat{\psi}^{\text{2S}}$ is asymptotically unbiased if at least one of the following conditions holds:.
  1. (i)
    $\|\hat{\pi}_{1}(X)-\pi_{1}(X)\|_{2}=o_{p}(1),$ and $\|\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\|_{2}=o_{p}(1)$;
  2. (ii)
    $\|\hat{\mu}_{1}(M,X)-\mu_{1}(M,X)\|_{2}=o_{p}(1),$ and $\|\hat{\mu}_{2}(X)-\mu_{2}(X)\|_{2}=o_{p}(1)$;
  3. (iii)
    $\|\hat{\pi}_{1}(X)-\pi_{1}(X)\|_{2}=o_{p}(1),$ and $\|\hat{\mu}_{1}(M,X)-\mu_{1}(M,X)\|_{2}=o_{p}(1)$.
Figure 2 summarizes the result in Proposition 6.
$\hat{\pi}_{1}$$\hat{\pi}_{2}$$\hat{\mu}_{1}$$\hat{\mu}_{2}$Stage 1:Stage 2:
Figure 2: Nuisance functions estimated in each stage of the approach. The arrow from $\hat{\pi}_{1}$ to $\hat{\pi}_{2}$ indicates that the estimator of $\pi_{1}$ is used in estimating $\pi_{2}$. The arrow from $\hat{\mu}_{1}$ to $\hat{\mu}_{2}$ is interpreted similarly. Correct specification of any of the pairs of functions in an ellipse suffices unbiasedness of the final estimator of the parameter of interest.
5.2 Asymptotic Normality
We next provide sufficient conditions for root-$n$ consistency and asymptotic normality of the estimator $\hat{\psi}^{\text{2S}}$.
Assumption 3.
  1. (a)
    $\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}=o_{p}(1)$, $\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2}=o_{p}(1)$, $\|\mu_{1}^{2}(M,X)-\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}=o_{p}(1)$, and $\|\mu_{2}^{2}(X)-\hat{\mu}_{2,\ell}^{2}(X)\|_{2}=o_{p}(1)$.
  2. (b)
    $\|(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\|_{2}=o_{p}(1)$ and $\|(\pi_{2}(M,X)-\hat{\pi}_{2,\ell}(M,X))^{2}\|_{2}=o_{p}(1)$.
  3. (c)
    There exists some constant $C$ such that $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $E[Y^{4}]$, are upper bounded by $C$.
  4. (d)
    $\big\|p(A=0\mid X)\hat{\pi}_{1,\ell}(X)-1\big\|_{2}\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2}=o_{p}(n^{-1/2})$, and
    $\big\|p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)\big\|_{2}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}=o_{p}(n^{-1/2})$.
For any choice of the nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, define,
$\Phi(O;\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2},\psi):=\phi(O;\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2})-\psi.$
We have the following result regarding the asymptotic behavior of $\hat{\psi}^{\text{2S}}$.
Theorem 2.
Under Assumptions 2 and  3, the estimator $\hat{\psi}^{\text{2S}}$ is $\sqrt{n}$-consistent and asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{\text{2S}}-\psi_{0})=\sqrt{n}E_{n}\left[\Phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2},\psi_{0})\right]+o_{p}(1).$
As a corollary of Theorem 2, one can use the influence function of $\psi_{0}$ to obtain confidence intervals for the parameter of interest. For $\ell\in\{1,\dots,L\}$, we estimate the variance of the influence function on fold $I_{\ell}$ as,
$\hat{\sigma}_{\ell}^{2}=E_{m}\big[\Phi^{2}(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell},\hat{\psi}^{\text{2S}})\big],$
and define,
$\hat{\sigma}^{2}=\frac{1}{L}\sum_{\ell=1}^{L}\hat{\sigma}_{\ell}^{2}.$
Then the $(1-\alpha)$ confidence interval of $\psi_{0}$ can be obtained as,
$\hat{\psi}^{\text{2S}}\pm z_{1-\alpha/2}\frac{\hat{\sigma}}{\sqrt{n}},$
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution.
6 Simulation Study
In this section, we evaluate our proposed method on synthetic data for estimating the mediation functional $\psi_{0}$. We consider a data generating process (DGP) which satisfies Assumptions 1 and 2, and study the robustness property of our estimator under the scenarios specified in Assumption 3. In particular, consider a three-dimensional covariate vector $X=(X_{1},X_{2},X_{3})$. The continuous covariates $X_{1}$ and $X_{2}$ are independently generated from the standard normal distribution $\mathcal{N}(0,1)$, and the binary covariate $X_{3}$ is generated from a Bernoulli distribution with $p(X_{3}=1)=0.5$. Treatment $A$ is generated according to
$p(A=1\mid X_{1},X_{2},X_{3})=\text{expit}(0.35+0.1X_{1}-0.2X_{2}+0.3X_{3}-0.15X_{1}^{2}+0.25X_{2}X_{3}).$
The mediator $M$ is drawn from a two-component Gaussian mixture centered around
$\mu_{M}(X_{1},X_{2},X_{3},A)=2.4+1.6X_{1}-1.2X_{2}+2.2X_{3}+0.8A+1.5X_{1}^{2}-1.8X_{2}X_{3}.$
Specifically,
$M\sim\begin{cases}\mathcal{N}(\mu_{M}(X_{1},X_{2},X_{3},A)-\delta,\sigma^{2}),&\text{with probability }p_{\text{left}},\\ \mathcal{N}(\mu_{M}(X_{1},X_{2},X_{3},A)+\delta,\sigma^{2}),&\text{with probability }1-p_{\text{left}},\end{cases}$
where $\sigma^{2}=1$ is the within-mode variance, $\delta=4$ is half the separation between the two modes, and $p_{\text{left}}=0.5$ controls the mixing proportion. Finally, the outcome $Y$ is generated from a Gaussian distribution with mean function
$\mu_{Y}(X_{1},X_{2},X_{3},M,A)$
$=6.4+2.7X_{1}+0.7X_{2}-3.6X_{3}-0.8X_{2}^{2}+2.5X_{1}A+3.2MA-2.9M+4.5A,$
and variance fixed at 1.
We consider sample sizes $n\in\{1000,2000,4000\}$ and repeat simulations under each sample size for 300 times. Cross-fitting is implemented with four folds, and hyperparameters for estimating $\pi_{1}$ and $\pi_{2}$ are tuned via cross-validation using a 75/25 train-validation split. For comparison, we also present the results for the TTS estimator $\hat{\psi}^{\text{TTS}}$ and a naive estimator $\hat{\psi}^{\text{naive}}=\frac{1}{L}\sum_{\ell=1}^{L}E_{m}[\hat{\mu}_{2,\ell}(X)]$, which is constructed from the identification formula $E\{E[E[Y\mid A=1,M,X]\mid A=0,X]\}$. In our implementation, $\mu_{1}$ and $\mu_{2}$ are estimated via random forests. The naive estimator uses the same $\mu_{2}$, while the TTS estimator also relies on random forests for nuisance estimation, except for the conditional density of the mediator, which is obtained using kernel density estimator. Replication code for the simulation study is available at https://github.com/changliues/TwoStage.git.
Table 1: Results for comparing TTS, naive, and two-stage estimators.
Sample Size Estimators Mean Bias Median Bias RMSE SD
n=1000 $\hat{\psi}^{\text{TTS}}$ $2.17\times 10^{33}$ $0.662\times 10^{-1}$ $3.70\times 10^{34}$ $3.70\times 10^{34}$
$\hat{\psi}^{\text{naive}}$ $1.59\times 10^{-1}$ $1.70\times 10^{-1}$ $2.69\times 10^{-1}$ $2.17\times 10^{-1}$
$\hat{\psi}^{\text{2S}}$ $1.68\times 10^{-1}$ $1.80\times 10^{-1}$ $2.76\times 10^{-1}$ $2.18\times 10^{-1}$
n=2000 $\hat{\psi}^{\text{TTS}}$ $4.15\times 10^{39}$ $2.47\times 10^{-1}$ $7.19\times 10^{40}$ $7.19\times 10^{40}$
$\hat{\psi}^{\text{naive}}$ $1.29\times 10^{-1}$ $1.40\times 10^{-1}$ $1.81\times 10^{-1}$ $1.27\times 10^{-1}$
$\hat{\psi}^{\text{2S}}$ $1.18\times 10^{-1}$ $1.22\times 10^{-1}$ $1.76\times 10^{-1}$ $1.31\times 10^{-1}$
n=4000 $\hat{\psi}^{\text{TTS}}$ $3.70\times 10^{60}$ $7.70\times 10^{-1}$ $1.07\times 10^{149}$ $1.07\times 10^{149}$
$\hat{\psi}^{\text{naive}}$ $13\times 10^{-1} 1.1$$16\times 10^{-1}$ $1.48\times 10^{-1}$ $0.965\times 10^{-1}$
$\hat{\psi}^{\text{2S}}$ $0.686\times 10^{-1}$ $0.696\times 10^{-1}$ $1.20\times 10^{-1}$ $0.986\times 10^{-1}$
Table 2: 95% confidence intervals and coverage for the two-stage estimator.
Sample Size 95% CI Coverage
n=1000 [9.590, 10.346] 0.870
n=2000 [9.643, 10.191] 0.883
n=4000 [9.670, 10.066] 0.910
Table 3: Robustness results for the two-stage estimator: Bias (RMSE).
Sample size all true $\{\pi_{1},\pi_{2}\}$ $\{\mu_{1},\mu_{2}\}$ $\{\pi_{1},\mu_{1}\}$ all false
n=1000 0.168 (0.276) 0.156 (0.276) 0.189 (0.287) 0.163 (0.273) 0.163 (0.273)
n=2000 0.118 (0.176) 0.137 (0.193) 0.184 (0.224) 0.155 (0.202) 0.185 (0.224)
n=4000 0.0686 (0.120) 0.108 (0.149) 0.160 (0.187) 0.140 (0.172) 0.203 (0.226)
Table 1 compares the TTS estimator, naive estimator, and two-stage estimator. The TTS estimator is highly unstable when the mediator is continuous, with mean bias and RMSE becoming excessively large due to the density in the denominator. In contrast, the naive estimator and the two-stage estimator remain stable. However, the proposed two-stage estimator exhibits lower bias and consistently outperforms the naive estimator in large samples. Results in Table 1 demonstrate the consistency of the two stage estimator. Moreover, as shown in Table 2, coverage for the two-stage estimator approaches the nominal 95% as sample size increases.
Table 3 illustrates the robustness property of the two-stage estimator. Each scenario corresponds to a setting in which specific subsets of nuisance functions are estimated well by the random forests, while the others are fit with suboptimal hyperparameters, matching the cases described in Proposition 6. Across the first four scenarios, the two-stage estimator remains stable, with bias decreasing as sample size grows, and it performs best when all nuisance functions are well-estimated. Even in the all false case, the estimator still remains stable. Together with Table 1, this confirms that the proposed two-stage estimator is both more accurate than the other two approaches and robust to nuisance function misspecification.
7 Application: Effect of Obesity on Cardiovascular Disease
In this section, we apply our proposed method to study the direct and indirect effect of obesity on coronary heart disease (CHD) mediated through the Glycohemoglobin (HbA1c). HbA1c, measured as a percentage of long-term blood glucose control, is the standard biomarker for diagnosing and monitoring diabetes, which is itself a major risk factor for CHD (Xu et al.,, 2017). Disentangling the extent to which obesity contributes to CHD risk through HbA1c, as opposed to other biological pathways, is of both scientific and policy relevance. If a substantial portion of the effect is mediated through HbA1c, interventions focused on glycemic control may effectively reduce obesity-related CHD risk; if not, alternative mechanisms such as inflammation, lipid metabolism, or hypertension may be more important to target. Partitioning the total effect into direct and indirect components thus provides critical insight for policymakers in determining whether diabetes prevention and management strategies are sufficient to mitigate CHD burden, or whether obesity itself must be targeted more aggressively regardless of glycemic control.
We use data from the 2013-2014 cycle of National Health and Nutrition Examination Survey (NHANES, National Center for Health Statistics, (2014)) for analysis. Obesity is defined using body mass index (BMI): individuals with BMI $\leq 29$ are classified as non-exposed (non-obese), and those with BMI $\geq 31$ are classified as exposed (obese). To sharpen the contrast between exposure groups, we exclude individuals with BMI in the range 29-31, leaving a deliberate gap around the threshold. The outcome of interest is a binary indicator based on self-reported physician diagnosis, where $Y=1$ indicates that the individual has previously been diagnosed with CHD. We adjust for eleven covariates that are potential confounders: age, gender, race, education level, annual family income, alcohol use, smoking history, vigorous work activity, diet quality, sleeping habits, and health insurance coverage. We perform complete case analysis, yielding a final sample size of 2685 and report effects in terms of odds ratios (OR) for ease of interpretation.
To motivate the definitions of direct and indirect ORs, recall from Section 2 that the ATE can be decomposed into the sum of the NIE and NDE. In an analogous manner, when considering binary outcomes, we define odds ratios in terms of the corresponding counterfactual event odds. The natural direct effect odds ratio is defined as
$\text{OR}_{\text{NDE}}=\frac{p(Y^{(1,M^{(0)})}=1)/(1-p(Y^{(1,M^{(0)})}=1))}{p(Y^{(0,M^{(0)})}=1)/(1-p(Y^{(0,M^{(0)})}=1))},$
the natural indirect effect odds ratio is defined as
$\text{OR}_{\text{NIE}}=\frac{p(Y^{(1,M^{(1)})}=1)/(1-p(Y^{(1,M^{(1)})}=1))}{p(Y^{(1,M^{(0)})}=1)/(1-p(Y^{(1,M^{(0)})}=1))},$
and the total effect odds ratio equals their product, $\text{OR}_{\text{TOT}}=\text{OR}_{\text{NIE}}\times\text{OR}_{\text{NDE}}$. While our earlier parameter of interest, $\theta_{0}=E[Y^{(1,M^{(0)})}]$, was expressed as a mean potential outcome, note that because the outcome is binary, this mean coincides with the probability $p(Y^{(1,M^{(0)})}=1)$, which then serves as the basis for constructing odds ratios. Since the denominator of $\text{OR}_{\text{NDE}}$ and numerator of $\text{OR}_{\text{NIE}}$ do not involve the mediator mechanism, we use the estimator $\frac{1}{L}\sum_{\ell=1}^{L}E_{m}[\hat{p}_{\ell}(Y=1\mid A=a,X)]$ for $a\in\{0,1\}$.
[…]
Taken together, our findings suggest that obesity results in approximately 60% higher odds of CHD compared to non-obesity, and that, considering only the direct pathway independent of HbA1c, obesity leads to about 35% higher odds of CHD.
8 Conclusion
When estimating a functional of interest that depends on nuisance functions, naively aiming for perfect fits for these nuisance functions can make the resulting estimator highly sensitive, as even small errors in the nuisance function estimators may be amplified, especially in structures involving ratios of density functions or scores in the denominator. Instead, it is important to investigate the role of nuisance functions in the bias structure of the estimator. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional, tailored to the bias structure of the influence function-based estimator. Our proposed nuisance estimators are based on the approach of balancing that can target the bias of the parameter of interest directly and also can lead to an estimator which is less sensitive to error. We demonstrated that the resulting estimator of the mediation function is multiply robust to the misspecification of the nuisance functions. Moreover, we provided conditions on the nuisance function estimators sufficient for obtaining root-$n$ consistency and asymptotic normality for the estimator of the mediation functional. We evaluated our proposed approach on synthetic data and applied our method to estimate the effect of obesity on coronary heart disease mediated by Glycohemoglobin, finding that obesity results in a 60% higher odds of coronary heart disease, of which roughly 35% is attributable to the direct pathway independent of Glycohemoglobin.
Appendix A Proofs
Proof of Proposition 2.
We note that
$\arg\max_{h_{1}}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}E\left[\left(\frac{1}{2}h_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}\right)^{2}\right]$
$=2E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg].$
Therefore,
$\breve{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}E\left[2\left\{(1-A)\pi_{1}(X)-1\right\}E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg]-E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg]^{2}\right]$
$=\arg\min_{\pi_{1}}E\left[E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg]^{2}\right],$
which since zero is achievable by $\frac{1}{p(A=0\mid X)}$, it implies that
$E\bigg[(1-A)\breve{\pi}_{1}(X)-1\bigg|X\bigg]=0.$
Therefore, for all $h_{1}\in\mathcal{H}_{1}$,
$E\left[\left\{(1-A)\breve{\pi}_{1}(X)-1\right\}h_{1}(X)\right]=E\left[E\bigg[(1-A)\breve{\pi}_{1}(X)-1\bigg|X\bigg]h_{1}(X)\right]=0.$
Proof of Proposition 3.
For any give $\hat{\pi}_{1}$, we note that
$\arg\max_{h_{2}}E\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{h_{2}}E\left[\left\{A\pi_{2}(M,X)(1-A)\hat{\pi}_{1}(X)\right\}^{2}+\frac{1}{4}h^{2}_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)\right]$
$=\arg\min_{h_{2}}E\left[\left(\frac{1}{2}h_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\right)^{2}\right]$
$=2E\bigg[A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg|M,X\bigg].$
Therefore,
$\breve{\pi}_{2}$ $=\arg\min_{\pi_{2}}\max_{h_{2}}E\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{\pi_{2}}E\bigg[2\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}E\big[A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\big|M,X\big]$
$~~~~~~~~~~~~~~~~~~~~-E\big[A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\big|M,X\big]^{2}\bigg]$
$=\arg\min_{\pi_{2}}E\left[E\bigg[A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg|M,X\bigg]^{2}\right],$
which since zero is achievable by $\frac{p(M,A=0\mid X)\hat{\pi}_{1}(X)}{p(M,A=1\mid X)}$, it implies that
$E\bigg[A\breve{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg|M,X\bigg]=0.$
Therefore, for all $h_{2}\in\mathcal{H}_{2}$,
$E\left[\left\{A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
$=E\left[E\bigg[A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\bigg|M,X\bigg]h_{2}(M,X)\right]=0.$
Proof of Proposition 4.
We first show that, for any function $\pi_{1}$, we have
$\max_{h_{1}\in\mathcal{H}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
where $\xi_{n}(\pi_{1})=\frac{1}{n}\big((1-A_{i})\pi_{1}(X_{i})-1\big)_{i=1}^{n}$.
To see this, by representer theorem, the solution to the maximization will be of the form,
$h_{1}(x)=\sum_{j=1}^{n}\alpha^{\ast}_{j}K_{\mathcal{H}_{1}}(X_{j},x),$
and it suffices to find the coefficients $\alpha^{\ast}=(\alpha_{j}^{\ast})_{j=1}^{n}$.
Note that we have,
$E_{n}[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)]$ $=\sum_{i=1}^{n}h_{1}(X_{i})\{\xi_{n}(\pi_{1})\}_{i}={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\},$
$E_{n}[h_{1}^{2}(X)]$ $=\frac{1}{n}{\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}^{2}{\alpha^{\ast}},$
$\|h_{1}\|_{\mathcal{H}_{1}}^{2}$ $={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}{\alpha^{\ast}}.$
Therefore we have,
$E_{n}\Big[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\}-{\alpha^{\ast}}^{\top}(\frac{1}{4n}K_{\mathcal{H}_{1},n}^{2}+\lambda_{\mathcal{H}_{1}}K_{\mathcal{H}_{1},n}){\alpha^{\ast}}.$
Taking the derivative and setting it to zero, the optimal coefficients are obtained as,
$\alpha^{\ast}=\frac{1}{2}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
and we have,
$\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\Big[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$=\frac{1}{4}\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\}.$
Therefore, the outer minimization problem is reduced to:
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\{\xi_{n}(\pi_{1})\}^{\top}\Gamma_{1}\{\xi_{n}(\pi_{1})\}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$
where $\Gamma_{1}=\frac{1}{4}K_{\mathcal{H}_{1},n}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$
By representer theorem, the solution to the maximization will be of the form,
$\pi_{1}(x)=\sum_{j=1}^{n}\alpha_{j}K_{\Pi_{1}}(X_{j},x),$
and it suffices to find the coefficients $\alpha=(\alpha_{j})_{j=1}^{n}$.
Note that,
$\xi_{n}(\pi_{1})$ $=\frac{1}{n}(diag(1-A)K_{\Pi_{1},n}\alpha-e_{n}),$
$\|\pi_{1}\|_{\Pi_{1}}^{2}$ $=\alpha^{\top}K_{\Pi_{1},n}\alpha.$
Therefore, we have,
$\min_{\pi_{1}\in\Pi_{1}}\{\xi_{n}(\pi_{1})\}^{\top}\Gamma_{1}\{\xi_{n}(\pi_{1})\}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2}$
$=\min_{\alpha\in\mathbb{R}^{n}}\frac{1}{n^{2}}\alpha^{\top}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha-\frac{2}{n^{2}}e_{n}^{\top}\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\lambda_{\Pi_{1}}\alpha^{\top}K_{\Pi_{1},n}\alpha+c,$
which is solved by
$\alpha=\Big(K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big)^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n}.$
The derivation for $\beta$ is similar and omitted for brevity.
Proof of Proposition 5.
  1. (a)
    For all $h_{1}\in L^{2}(P_{0,X})$, we have
    $E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=E\left[\left\{E[1-A\mid X]\pi_{1}(X)-1\right\}h_{1}(X)\right]=0.$
  2. (b)
    For all $h_{2}\in L^{2}(P_{0,(M,X)})$, we have
    $E\left[\left\{A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
    $=E\left[\left\{\frac{p(A=1\mid M,X)p(M\mid A=0,X)}{p(M,A=1\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=E\left[\left\{\frac{p(M\mid A=0,X)}{p(M\mid X)}-\frac{p(A=0\mid M,X)}{p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=E\left[\left\{\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}-\frac{p(M,A=0\mid X)}{p(M\mid X)p(A=0\mid X)}\right\}h_{2}(M,X)\right]$
    $=0.$
Proof of Proposition 6.
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$
$=E[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}]-\psi_{0}$
$=E\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~~~+E\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[\left\{E\left[1-A\mid X\right]\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~~~+E\left[\left\{E\left[A\mid M,X\right]\hat{\pi}_{2}(M,X)-E\left[1-A\mid M,X\right]\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~~~+E\left[\left\{p\left(A=1\mid M,X\right)\hat{\pi}_{2}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
Using Cauchy-Schwarz inequality, we have
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$
$\leq\Big\|p\left(A=0\mid X\right)\hat{\pi}_{1}(X)-1\Big\|_{2}\Big\|\mu_{2}(X)-\hat{\mu}_{2}(X)\Big\|_{2}$
$~~~~~+\Big\|p\left(A=1\mid M,X\right)\hat{\pi}_{2}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1}(X)\Big\|_{2}\Big\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big\|_{2}$
$\leq\Big\|p\left(A=0\mid X\right)\Big(\hat{\pi}_{1}(X)-\pi_{1}(X)\Big)\Big\|_{2}\Big\|\mu_{2}(X)-\hat{\mu}_{2}(X)\Big\|_{2}$
$~~~~~+\Big\|\frac{p\left(A=1\mid M,X\right)}{\pi_{1}(X)}\Big(\pi_{1}(X)\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\hat{\pi}_{1}(X)\Big)\Big\|_{2}\Big\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big\|_{2}$
$\leq\Big\|\hat{\pi}_{1}(X)-\pi_{1}(X)\Big\|_{2}\Big\|\mu_{2}(X)-\hat{\mu}_{2}(X)\Big\|_{2}$
$~~~~~+\Big(\Big\|\pi_{1}(X)\Big(\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\Big)\Big\|_{2}+\Big\|\pi_{2}(M,X)\Big(\hat{\pi}_{1}(X)-\pi_{1}(X)\Big)\Big\|_{2}\Big)\Big\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big\|_{2}$
$\leq\Big\|\hat{\pi}_{1}(X)-\pi_{1}(X)\Big\|_{2}\Big\|\mu_{2}(X)-\hat{\mu}_{2}(X)\Big\|_{2}$
$~~~~~+\frac{1}{\epsilon^{2}}\Big(\Big\|\hat{\pi}_{2}(M,X)-\pi_{2}(M,X)\Big\|_{2}+\Big\|\hat{\pi}_{1}(X)-\pi_{1}(X)\Big\|_{2}\Big)\Big\|\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\Big\|_{2},$
where the last equation holds by Assumption 2. Therefore if at least one of the conditions in the proposition holds, we have
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$ $=o_{p}(1).$
Proof of Theorem 1.
$E[\hat{\psi}^{\text{2S}}]-\psi_{0}$
$=E\left[\hat{\mu}_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\right]-E[\mu_{2}(X)]$
$=E\left[\hat{\mu}_{2}(X)-\mu_{2}(X)+(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[E\left[(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{2}(X)\}\bigg|X\right]-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[E\left[A\hat{\pi}_{2}(M,X)\{Y-\hat{\mu}_{1}(M,X)\}\bigg|M,X\right]-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[E[(1-A)\mu_{1}(M,X)\mid X]\hat{\pi}_{1}(X)-\frac{\hat{\pi}_{1}(X)\hat{\mu}_{2}(X)}{\pi_{1}(X)}-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\Big[\hat{\pi}_{2}(M,X)\big(E[AY\mid M,X]-E[A\mid M,X]\hat{\mu}_{1}(M,X)\big)-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\Big]$
$=E\left[E[(1-A)\mid X]\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\hat{\pi}_{1}(X)-\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[\hat{\pi}_{2}(M,X)E[A\mid M,X]\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}-(1-A)\hat{\pi}_{1}(X)\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right]$
$=E\left[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2}(X)\}\right]$
$~~~+E\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\}\right].$
Proof of Theorem 2.
Assume the size of each fold of the data is $m$, that is, $n=mL$. Let $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ be the estimators of $(\pi_{1},\pi_{2},\mu_{1},\mu_{2})$ obtained on subsamples $I_{\ell}^{c}$. Then $\hat{\psi}^{\text{2S}}_{\ell}$ can be written as
$\hat{\psi}^{\text{2S}}_{\ell}=E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right].$
Therefore, we have
$\sqrt{n}(\hat{\psi}^{\text{2S}}-\psi_{0})$ $=\sqrt{n}\Big\{\frac{1}{L}\sum_{\ell=1}^{L}E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}$
$=\frac{\sqrt{n}}{L\sqrt{m}}\sum_{\ell=1}^{L}\Bigg\{\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\}$
$\quad\quad\quad\quad\quad\quad\quad+\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$
$\quad\quad\quad\quad\quad\quad\quad+\sqrt{m}\Big\{E\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}\Bigg\},$
where for each function $f(O)$,
$\mathbb{G}_{n}[f]=\sqrt{n}\{E_{n}[f]-E[f]\},$
represents the empirical process. Recalling that $n=mL$, we have
$\sqrt{n}(\hat{\psi}^{\text{2S}}-\psi_{0})$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\}$ $(T_{1})$
$\quad+\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$ $(T_{2})$
$\quad+\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big\{E\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}.$ $(T_{3})$
We will show that $(T_{2})$ provides us with the term $\sqrt{n}E_{n}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$, and under the assumptions of the theorem, $(T_{1})$ and $(T_{3})$ are $o_{p}(1)$.
Analysis of $T_{1}$
Define
$A_{m}^{\ell}:=\left\{\mathbb{G}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\right\},$
where the empirical processes are evaluated on $I_{\ell}$. We note that
$var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)$
$=m\cdot var\left(E_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-E_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]\mid I_{\ell}^{c}\right)$
$=var\big(\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})-\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\mid I_{\ell}^{c}\big)$
$\leq E\big[\{\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})-\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\}^{2}\mid I_{\ell}^{c}\big]$
$=E\Big[\Big\{\hat{\mu}_{2,\ell}(X)+(1-A)\hat{\pi}_{1,\ell}(X)\{\hat{\mu}_{1,\ell}(M,X)-\hat{\mu}_{2,\ell}(X)\}+A\hat{\pi}_{2,\ell}(M,X)\{Y-\hat{\mu}_{1,\ell}(M,X)\}$
$\quad\quad-\mu_{2}(X)-(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}-A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big\}^{2}\Big|I_{\ell}^{c}\Big]$
$=E\Big[\Big\{\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
$\quad\quad-\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)-\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)-AY\pi_{2}(M,X)\Big\}^{2}\Big|I_{\ell}^{c}\Big]$
$\leq 5\Big(\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|(1-A)(\hat{\mu}_{2,\ell}(X)\hat{\pi}_{1,\ell}(X)-\mu_{2}(X)\pi_{1}(X))\|_{2}^{2}$
$\quad\quad+\|(1-A)(\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{1,\ell}(X)-\mu_{1}(M,X)\pi_{1}(X))\|_{2}^{2}$
$\quad\quad+\|A(\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{2,\ell}(M,X)-\mu_{1}(M,X)\pi_{2}(M,X))\|_{2}^{2}+\|AY(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}\Big)$
$\leq 5\Big(\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|\hat{\mu}_{2,\ell}(X)\hat{\pi}_{1,\ell}(X)-\mu_{2}(X)\pi_{1}(X)\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{1,\ell}(X)-\mu_{1}(M,X)\pi_{1}(X)\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)\hat{\pi}_{2,\ell}(M,X)-\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}^{2}+\|Y(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}\Big)$
$\leq 10\Big(\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|\hat{\mu}_{2,\ell}(X)(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))\|_{2}^{2}+\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))\|_{2}^{2}+\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}(M,X)(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}+\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|Y(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\|_{2}^{2}\Big)$
$\leq 10\Big(\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}^{2}+\|\hat{\mu}_{2,\ell}^{2}(X)\|_{2}\|(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))^{2}\|_{2}+\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}\|(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))^{2}\|_{2}+\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\|_{2}+\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}^{2}$
$\quad\quad+\|Y^{2}\|_{2}\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\|_{2}\Big).$
By Assumption 3 (a) and (b), $\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}$, $\|\hat{\mu}_{2,\ell}(X)-\mu_{2}(X)\|_{2}$, $\|\hat{\mu}_{1,\ell}^{2}(M,X)-\mu_{1}^{2}(M,X)\|_{2}$, $\|\hat{\mu}_{2,\ell}^{2}(X)-\mu_{2}^{2}(X)\|_{2}$, $\|(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X))^{2}\|_{2}$, and $\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\|_{2}$ converge to zero in probability. By Assumption 2,
$\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}$ $\leq\frac{1}{\epsilon}\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\|_{2},$
$\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$ $\leq\frac{1}{\epsilon}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2},$
$\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$ $=\|\frac{p(A=0\mid M,X)}{p(A=1\mid M,X)}\pi_{1}(X)(\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X))\|_{2}$
$\leq\frac{1}{\epsilon^{2}}\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\|_{2}.$
Therefore, $\|\pi_{1}(X)(\hat{\mu}_{2,\ell}(X)-\mu_{2}(X))\|_{2}$, $\|\pi_{1}(X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$, and $\|\pi_{2}(M,X)(\hat{\mu}_{1,\ell}(M,X)-\mu_{1}(M,X))\|_{2}$ also converge to zero in probability.
By Assumption 3 (c), $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $\|Y^{2}\|_{2}$ are bounded by some constants, which imply that
$\|\hat{\mu}_{1,\ell}^{2}(M,X)\|_{2}$ $\leq\|\hat{\mu}_{1,\ell}^{2}(M,X)-\mu_{1}^{2}(M,X)\|_{2}+\|\mu_{1}^{2}(M,X)\|_{2}\leq o_{p}(1)+C=O_{p}(1),$ (7)
$\|\hat{\mu}_{2,\ell}^{2}(X)\|_{2}$ $\leq\|\hat{\mu}_{2,\ell}^{2}(X)-\mu_{2}^{2}(X)\|_{2}+\|\mu_{2}^{2}(X)\|_{2}\leq o_{p}(1)+C=O_{p}(1).$ (8)
Therefore,
$var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)=o_{p}(1).$
Also note that,
$E\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]$
$=E\Big[\hat{\mu}_{2,\ell}(X)+(1-A)\hat{\pi}_{1,\ell}(X)\{\hat{\mu}_{1,\ell}(M,X)-\hat{\mu}_{2,\ell}(X)\}+A\hat{\pi}_{2,\ell}(M,X)\{Y-\hat{\mu}_{1,\ell}(M,X)\}$
$\quad\quad-\mu_{2}(X)-(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}-A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big|I_{\ell}^{c}\Big]$
$=E\Big[\hat{\mu}_{2,\ell}(X)\left(1-(1-A)\hat{\pi}_{1,\ell}(X)\right)+\hat{\mu}_{1,\ell}(M,X)\left((1-A)\hat{\pi}_{1,\ell}(X)-A\hat{\pi}_{2,\ell}(M,X)\right)+AY\hat{\pi}_{2,\ell}(M,X)$
$\quad\quad-\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)-\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)-AY\pi_{2}(M,X)\Big|I_{\ell}^{c}\Big]$
$=E\Big[\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)\left(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)\right)$
$\quad\quad-\mu_{2}(X)\underbrace{\left(1-p(A=0\mid X)\pi_{1}(X)\right)}_{=0}$
$\quad\quad-\mu_{1}(M,X)\underbrace{\left(p(A=0\mid M,X)\pi_{1}(X)-p(A=1\mid M,X)\pi_{2}(M,X)\right)}_{=0}$
$\quad\quad+\mathbb{E}[AY\mid M,X]\hat{\pi}_{2,\ell}(M,X)-\mathbb{E}[AY\mid M,X]\pi_{2}(M,X)\Big|I_{\ell}^{c}\Big]$
$=E\Big[\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)\left(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)\right)$
$\quad\quad+p(A=1\mid M,X)\mu_{1}(M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=1\mid M,X)\mu_{1}(M,X)\pi_{2}(M,X)\Big|I_{\ell}^{c}\Big].$
By Jensen’s inequality and Cauchy-Schwarz inequality,
$E\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]$
$\leq\Big\|\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)\left(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X)\right)$
$\quad\quad+p(A=1\mid M,X)\mu_{1}(M,X)\hat{\pi}_{2,\ell}(M,X)-p(A=1\mid M,X)\mu_{1}(M,X)\pi_{2}(M,X)\Big\|_{2}$
$=\Big\|\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)$
$\quad\quad+\hat{\mu}_{1,\ell}(M,X)(p(A=0\mid M,X)\hat{\pi}_{1,\ell}(X)-p(A=0\mid M,X)\pi_{1}(X)$
$\quad\quad\quad+p(A=1\mid M,X)\pi_{2}(M,X)-p(A=1\mid M,X)\hat{\pi}_{2,\ell}(M,X))$
$\quad\quad+p(A=1\mid M,X)\mu_{1}(M,X)(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\Big\|_{2}$
$\leq\Big\|\hat{\mu}_{2,\ell}(X)\left(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X)\right)\Big\|_{2}$
$\quad\quad+\Big\|\hat{\mu}_{1,\ell}(M,X)p(A=0\mid M,X)\left(\hat{\pi}_{1,\ell}(X)-\pi_{1,\ell}(X)\right)\Big\|_{2}$
$\quad\quad+\Big\|\mu_{1,\ell}(M,X)p(A=1\mid M,X)(\pi_{2}(M,X)-\hat{\pi}_{2,\ell}(M,X))\Big\|_{2}$
$\quad\quad+\Big\|p(A=1\mid M,X)\mu_{1}(M,X)(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))\Big\|_{2}$
$\leq\Big\|\hat{\mu}_{2,\ell}^{2}(X)\Big\|_{2}^{1/2}\Big\|(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}^{1/2}$
$\quad\quad+\Big\|(p(A=0\mid M,X)\hat{\mu}_{1,\ell}(M,X))^{2}\Big\|_{2}^{1/2}\Big\|\big(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X)\big)^{2}\Big\|_{2}^{1/2}$
$\quad\quad+\Big\|(p(A=1\mid M,X)\hat{\mu}_{1,\ell}(M,X))^{2}\Big\|_{2}^{1/2}\Big\|\big(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X)\big)^{2}\Big\|_{2}^{1/2}$
$\quad\quad+\Big\|(p(A=1\mid M,X)\mu_{1}(M,X))^{2}\Big\|_{2}^{1/2}\Big\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\Big\|_{2}^{1/2}$
$\leq\Big\|\hat{\mu}_{2,\ell}^{2}(X)\Big\|_{2}^{1/2}\Big\|(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}^{1/2}+\Big\|\hat{\mu}_{1,\ell}^{2}(M,X)\Big\|_{2}^{1/2}\Big\|\big(\hat{\pi}_{1,\ell}(X)-\pi_{1}(X)\big)^{2}\Big\|_{2}^{1/2}$
$\quad\quad+\Big\|\hat{\mu}_{1,\ell}^{2}(M,X)\Big\|_{2}^{1/2}\Big\|\big(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X)\big)^{2}\Big\|_{2}^{1/2}+\Big\|\mu_{1}^{2}(M,X)\Big\|_{2}^{1/2}\Big\|(\hat{\pi}_{2,\ell}(M,X)-\pi_{2}(M,X))^{2}\Big\|_{2}^{1/2}.$
Also note that,
$\Big\|(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}$ $=\Big\|(p(A=0\mid X))^{2}(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}$
$\leq\Big\|(\pi_{1}(X)-\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}.$
Hence, by Assumption 3 (b), $\Big\|(1-p(A=0\mid X)\hat{\pi}_{1,\ell}(X))^{2}\Big\|_{2}$ converges to zero in probability.
Similar to inequalities (7) and (8), $\Big\|\hat{\mu}_{1,\ell}^{2}(M,X)\Big\|_{2}^{1/2}$ and $\Big\|\hat{\mu}_{2,\ell}^{2}(X)\Big\|_{2}^{1/2}$ are $O_{p}(1)$. Therefore, by Assumption 3 (a), (b) and (c),
$E\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]=o_{p}(1).$
Therefore,
$E\left[(A_{m}^{\ell})^{2}\mid I_{\ell}^{c}\right]=var\left(A_{m}^{\ell}\mid I_{\ell}^{c}\right)+\Big(E\left[A_{m}^{\ell}\mid I_{\ell}^{c}\right]\Big)^{2}\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
Then by conditional Chebyshev inequality, for all $\delta>0$, we have,
$P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
That is, $P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})$ is a bounded sequence that converges to zero in probability. Therefore,
$E\left[P(|A_{m}^{\ell}|>\delta\mid I_{\ell}^{c})\right]=P(|A_{m}^{\ell}|>\delta)\stackrel{{\scriptstyle p.}}{{\rightarrow}}0,\quad\text{ as }m\rightarrow\infty.$
That is, $A_{m}^{\ell}\stackrel{{\scriptstyle p.}}{{\rightarrow}}0$ as $m\rightarrow\infty$.
The conclusion holds for all $\ell\in\{1,\dots,L\}$ and hence, we conclude that $T_{1}=o_{p}(1)$.
Analysis of $T_{2}$
$T_{2}$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\mathbb{G}_{m}\left[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})\right]$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{E_{m}[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]-E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\frac{1}{m}\sum_{\ell=1}^{L}\sum_{i=1}^{m}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\{\phi(O_{i};\pi_{1},\pi_{2},\mu_{1},\mu_{2})-E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})]\right\}.$
By central limit theorem, the last expression converges to the distribution $\mathcal{N}(0,var(\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2}))$ if $E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$. Note that $E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$ can be upper bounded as
$E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]$
$=E\Big[\Big\{\mu_{2}(X)+(1-A)\pi_{1}(X)\{\mu_{1}(M,X)-\mu_{2}(X)\}+A\pi_{2}(M,X)\{Y-\mu_{1}(M,X)\}\Big\}^{2}\Big]$
$=E\Big[\Big\{\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)+\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)+AY\pi_{2}(M,X)\Big\}^{2}\Big]$
$=\|\mu_{2}(X)\left(1-(1-A)\pi_{1}(X)\right)+\mu_{1}(M,X)\left((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\right)+AY\pi_{2}(M,X)\|_{2}^{2}$
$\leq\Big(\|\mu_{2}(X)\big(1-(1-A)\pi_{1}(X)\big)\|_{2}+\|\mu_{1}(M,X)\big((1-A)\pi_{1}(X)-A\pi_{2}(M,X)\big)\|_{2}$
$\quad\quad+\|AY\pi_{2}(M,X)\|_{2}\Big)^{2}$
$\leq\Big(\|\mu_{2}(X)\|_{2}+\|(1-A)\mu_{2}(X)\pi_{1}(X)\|_{2}+\|(1-A)\mu_{1}(M,X)\pi_{1}(X)\|_{2}+\|A\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}$
$\quad\quad+\|AY\pi_{2}(M,X)\|_{2}\Big)^{2}$
$\leq\Big(\|\mu_{2}(X)\|_{2}+\|\mu_{2}(X)\pi_{1}(X)\|_{2}+\|\mu_{1}(M,X)\pi_{1}(X)\|_{2}+\|\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}+\|Y\pi_{2}(M,X)\|_{2}\Big)^{2}$
By Assumption 2 and Jensen’s inequality,
$\|\mu_{2}(X)\|_{2}\leq\|\mu_{2}^{2}(X)\|_{2}^{1/2},$
$\|\mu_{2}(X)\pi_{1}(X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{2}(X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{2}^{2}(X)\|_{2}^{1/2},$
$\|\mu_{1}(M,X)\pi_{1}(X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{1}(M,X)\|_{2}\leq\frac{1}{\epsilon}\|\mu_{1}^{2}(M,X)\|_{2}^{1/2},$
$\|\mu_{1}(M,X)\pi_{2}(M,X)\|_{2}\leq\frac{1}{\epsilon^{2}}\|\mu_{1}(M,X)\|_{2}\leq\frac{1}{\epsilon^{2}}\|\mu_{1}^{2}(M,X)\|_{2}^{1/2},$
$\|Y\pi_{2}(M,X)\|_{2}\leq\frac{1}{\epsilon^{2}}\|Y\|_{2}\leq\frac{1}{\epsilon^{2}}\|Y^{2}\|_{2}^{1/2}.$
Therefore, by Assumption 3 (c), we have $E[\phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2})^{2}]<\infty$.
Analysis of $T_{3}$
Using result of Theorem 1 and Law of Total Probability, we have,
$T_{3}$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big\{E\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{E\left[\hat{\psi}^{\text{2S}}_{\ell}\right]-\psi_{0}\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{E\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~~~~~~+E\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{E\left[\left\{E\left[1-A\mid X\right]\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~~~~~~+E\left[\left\{E\left[A\mid M,X\right]\hat{\pi}_{2,\ell}(M,X)-E\left[1-A\mid M,X\right]\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{E\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~~~~~~~~~~~~~~~+E[\{p\left(A=1\mid M,X\right)\hat{\pi}_{2,\ell}(M,X)$
$~~~~~~~~~~~~~~~-p\left(A=0\mid M,X\right)\hat{\pi}_{1,\ell}(X)\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}]\Bigg\}.$
Using Cauchy-Schwarz inequality, we have
$T_{3}$ $\leq\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{\Big\|p\left(A=0\mid X\right)\hat{\pi}_{1,\ell}(X)-1\Big\|_{2}\Big\|\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\Big\|_{2}$
$~~~~~~+\Big\|p\left(A=1\mid M,X\right)\hat{\pi}_{2,\ell}(M,X)-p\left(A=0\mid M,X\right)\hat{\pi}_{1,\ell}(X)\Big\|_{2}\Big\|\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\Big\|_{2}\Bigg\}.$
Therefore by Assumption 3 (d), we have
$T_{3}$ $=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{o_{p}(n^{-1/2})+o_{p}(n^{-1/2})\Bigg\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\times L\times o_{p}(n^{-1/2})$
$=\sqrt{mL}\times o_{p}(n^{-1/2})$
$=\sqrt{n}o_{p}(n^{-1/2})$
$=o_{p}(1).$